### v1

# Chapter 0 Hadron Production Processes

Abstract
[Abstract]The experimental search for the pion – proposed in 1935 by Hideki Yukawa as the force carrier of the strong nucleon-nucleon interaction – was rewarded in 1947 when in cosmic ray photographic emulsion data a charged particle was identified with the proper mass of about 300 times the electron mass, completed three years later by the discovery of the neutral pion. Since then, accelerator-driven pion and meson (photo-)production on the nucleon and the associated production of new baryons have become the key elements for ground-breaking discoveries in numerous areas of particle and nuclear physics, from fundamental symmetries and their breaking to low-energy QCD dynamics, laying also foundations for modern elementary particle physics and the Standard Model. This article is an overview of eight decades of experimental and theoretical meson production physics, from isospin to charm and beyond, forming our understanding of hadrons and their interactions.
keywords:
[…]
Figure 1: Feynman diagram illustrating the production of a pion on a proton by an incident photon $\gamma$.
Nomenclature] AGS Alternating Gradient Synchrotron at BNL ANL Argonne National Laboratory ALICE A Large Ion Collider Experiment at CERN Bevatron Billions of eV synchrotron at LBL BEVALAC Bevatron plus HILAC linear accelerator at LBL BIC Bound States in the Continuum BNL Brookhaven National Laboratory BnGa Bonn-Gatchina Model BSE Bethe-Salpeter Equation CERN European Organization for Nuclear Research (Conseil Européen pour la Recherche Nucléaire) CC Model Coupled Channel Model CGLN amplitudes Chew, Goldberger, Low, and Nambu amplitudes CLAS CEBAF Large Acceptance Spectrometer at JLab c.m. Center-of-Mass Frame COSY Cooler Synchrotron CQM Constituent Quark Model ELSA Elektronen-Stretcher-Anlage (Electron-Stretcher-Facility) EM Electromagnetic FAIR Facility for Antiproton and Ion Research GANIL Grand Accélérateur National d’Ions Lourds GiM Giessen Model GlueX Gluonic Excitation Experiment GMO Gell-Mann-Okubo mass formula GRAAL Grenoble Accelerator LASer GSI Gesellschaft für Schwerionenforschung GW The George Washington University HADES High Acceptance DiElectron Spectrometer HILAC Heavy Ion Linear Accelerator at LBL ITEP Institute Theoretical and Experimental Physics JLab Jefferson Laboratory or Thomas Jefferson National Accelerator Facility (TJNAF) J-PARC Japan Proton Accelerator Research Complex LAMPF Los Alamos Meson Physics Facility LBL Lawrence Berkeley Laboratory LEPS Laser Electron Photon Experiment LET Low Energy Theorem LHC Large Hadron Collider at CERN LQCD Lattice Quantum Chromo-Dynamics MAMI Mainz Microtron OZI Rule Okubo–Zweig–Iizuka rule PDG Particle Data Group PSI Paul Scherrer Institute, previously SIN - Swiss Institute for Nuclear Research PWA Partial-Wave Analysis QCD Quantum Chromo-Dynamics QFT Quantum Field Theory QGP Quark–Gluon Plasma RICH Relativistic Heavy Ion Collider at BNL RIKEN National Research and Development Agency, Japan SAID Scattering Analyses Interactive Data SM Standard Model TRIUMF Tri-University Meson Facility VMD Model Vector Meson Dominance Model
Objectives
  • Short story of hadron physics.
  • Production of short-lived particles underlying strong interactions.
  • Spectrum of hadrons as asymptotic mass eigenstates of QCD.
  • Probing the structure and dynamics of hadrons.
  • Experimental methods and facilities.
  • Theoretical concepts and results.
1 Introduction
1 Emergence of Subnuclear Particle Physics
The first recorded observation of particles different from the material forming the everyday-terrestrial environment was made as early as 1912 by Victor Franz Hess (1883-1964) in high-altitude balloon campaigns [2]. First considered merely as a curiosity, soon after the wide-reaching importance of Hess’ observations became evident. For his groundbreaking research, Hess was awarded the Nobel Prize in Physics in 1936, together with Carl David Anderson (1905-1991), who discovered the positron in cosmic rays [3], thereby opening the era of antimatter research. The modern understanding of air showers produced by cosmic rays is illustrated in Fig. 2.
The epochal scientific revolution of quantum physics, speeding up in the decade after 1920, was necessary to make cosmic-ray research a flourishing enterprise. A big step forward was made by Patrick Maynard Stuart Blackett (1897-1974) by constructing in collaboration with Giuseppe Paolo Stanislao “Beppo” Occhialini (1907-1993) a coincidence Wilson cloud chamber. In 1933, he could confirm the positron and, in the years after, charged particle showers and pair production and pair annihilation processes, which he explained on the basis of the just formulated Dirac theory [^1] Paul Adrien Maurice Dirac (1902-1984) shared the 1933 Nobel Prize in Physics with Erwin Schrödinger (1887-1961) for the discovery of new productive forms of atomic theory. Cloud chambers, operated by supersaturated water vapor, were invented much earlier by the Scottish physicist Charles Thomson Rees Wilson (1869-1959), who was awarded the Nobel Prize in Physics in 1927 for his method of making the paths of electrically charged particles visible by condensation of vapor.
In the late 1930s, Cecil Frank Powell (1903-1969) joined cosmic-ray research. He and Occhialini had optimized the technology and the use of photographic plates, which became essential for cosmic-ray research - and for many years after. Exposing the plates at high altitude to cosmic rays, Lattes, Muirhead, Occhialini, and Powell recorded for the first time the decay of a pi-meson, $\pi^{-}\to\mu^{-}+\bar{\nu}_{\mu}$, shown in Fig. 3 and only later understood to be caused by weak interaction [4, 5]. This measurement can be considered as the start of research on hadron production, first evidenced by the track recording the existence of the decaying particle and second by identifying that particle and its properties by the decay products. Basically, the same scheme is applied in modern production experiments at accelerator facilities, albeit on a much more involved technological level. In 1950, Powell and Occhialini won the Nobel Prize for their seminal work.
In their first paper on the pion observation of 1947 [4], Lattes, Occhialini, and Powell write: We represent the primary mesons by the symbol $\pi$, and the secondary by $\mu$. In the second paper of 1947 [5], they call the particles casually pion and muon. There is good reason to state that this is the only documented birth and baptism certificate of the pion. The actual inventor of - at that time - such a hypothetical particle, however, was Hideki Yukawa (1907-1981) [6, 7] who postulated the existence of a meson (lent from the ancient Greek language for medium, i.e., between electron and nucleon masses). By that hypothesis, Yukawa could explain the short-ranged forces between nucleons in accordance with the extremely small size of nuclei. After the discovery of a medium-mass particle of about the required mass of about 200 $m_{e}$ in cosmic-ray showers, the new particle was initially considered as the Yukawa meson. By the work of Lattes et al., it became clear that the earlier found particle was not the Yukawa meson but the pion was the proper candidate while the muon was of a different nature. For a short while, the muon was in fact considered as the light meson but soon the distinct nature of the muon as a lepton and fermion [^2] Half-integer spin-particles are named after Enrico Fermi (1901-1954) who received the Noble price of 1938 for his demonstrations of the existence of new radioactive elements produced by neutron irradiation, and for his related discovery of nuclear reactions brought about by slow neutrons was understood.
Interestingly, Yukawa himself apparently never claimed openly the name pion or pi-meson of his origin. In his original papers of 1935 and 1937, Yukawa denotes the new quantum as U-field [6, 7]. In his Nobel lecture of 1949 [8] and in a paper of that year [9] Yukawa seems to have accepted the naming of his meson as pion or pi-meson by using those names repeatedly.
Figure 2: Illustration of an air shower resulting from the interaction of an incoming high energy cosmic particle with atmospheric nuclei. The figure is adapted from Ref. [10].
In his own words, Yukawa summarizes in [9] the situation of 1949 by: This rather puzzling situation of the meson theory changed in 1947 due to the experiment by the Rome group on the decay of negative mesons on the one hand, and the discovery by the Bristol group of two kinds of mesons in cosmic rays on the other., where the Bristol group are Lattes et al. and the Rome group refers to an earlier paper of 1943 by Nereson and Rossi [11]. By denoting the muon as negative meson, Yukawa continues to use the notation which had been established after about 1935 when it was thought that there would be only a single meson.
Figure 3: The photographic emulsion plate showing the first observation of - as known today - the weak decay of a pion, $\pi^{-}\to\mu^{-}+\bar{\nu}_{\mu}$. The muon-antineutrino $\bar{\nu}_{\mu}$ is not observed. Photo, including the original caption, taken from Ref. [4].
2 Hadron Production on Accelerators
As early as 1948, in parallel to cosmic ray research, new opportunities for hadron physics showed up when Gardner and Lattes published the first results on the Production of Mesons by the 184-Inch Berkeley Cyclotron [12], demonstrating that hadron research could be performed under laboratory conditions. The circulated beam of $380~{}\mathrm{MeV}$ alpha particles inside the cyclotron passed through the thin carbon target, producing mesons and a bunch of other particles. The negatively charged mesons were sorted out by the magnetic field and roughly focused on the edge of a stack of photographic plates. Besides using carbon targets, a few experiments were made on beryllium, copper, and uranium targets.
The modern era of particle physics - lasting until today - started with the advent of particle accelerators being able to deliver intense beams of highly energetic particles. In 1953, the Cosmotron synchrotron at BNL delivered proton beams with the design energy of $3.3~{}\mathrm{GeV}$. A year later, the Bevatron at Berkeley laboratory became operative with proton beams of even $6.2~{}\mathrm{GeV}$. With these machines, the cosmic ray mesons and V-particles could be produced and studied at the laboratory. The most eminent Bevatron result was - and is - the discovery of hadronic antimatter by observing first the antiproton in 1955 [13] and shortly after also the antineutron in 1956 [14], for which Owen Chamberlain (1920-1959) and Emilio Gino Segré (1905-1989) were honored with the Nobel Prize in Physics of 1959 for the discovery of the antiproton, a sub-atomic antiparticle.
1 Entering Strange Territory and Symmetry Violation
The years around 1947 actually played a key role in hadron discoveries. Rochester and Butler reported on Wilson cloud chamber observations indicating Evidence for the Existence of New Unstable Elementary Particles [15] showing an unusual decay pattern resembling a two-pronged fork, later denoted as V-particles [16], of mass $980~{}m_{e}\leq m<m_{p}$ ($m_{e}\simeq 511~{}\mathrm{keV}$ and $m_{p}\simeq 938.3~{}\mathrm{MeV}$ are the electron and proton rest masses, respectively) - fitting well with what shortly after was named the $K$-meson. In 1950, Hopper and Biswas of Melbourne observed another neutral particle decaying by the same V-like pattern seen by Rochester and Butler, but of a much heavier mass of about 2370 $m_{e}$ which matches surprisingly well the rest mass of the $\Lambda$ hyperon. The characteristic decay pattern of these strange particles led to the name V-particles [17]. Hence, the first hadrons - a meson and a baryon - with strangeness content were observed, although it still took another 15 years before a concise theoretical scheme was formulated.
In the years after, the strangeness sector could be explored systematically, resulting in a whole spectrum of higher lying hyperon states. The observation of the non-conservation of parity (P) in hyperon decays was the first experimental datum of a broken symmetry at the level of elementary particles. In many cases, however, detailed spectroscopic studies were beyond the technological capabilities of that time. A typical case is the $\Lambda(1520)$ excitation. That $J^{\pi}=\frac{3}{2}^{-}$ excited state of the $\Lambda$ hyperon was observed originally as an elastic resonance in kaon scattering on a proton target in 1962 [18]. However, it lasted about 50 years and required an electroproduction experiment and the advanced detector setup at JLab before the pole position for the first hyperon of that rather sharp resonance ($\Gamma\simeq 18~{}\mathrm{MeV}$) could be determined precisely for $\Lambda(1520)$  [19].
A far-reaching result of the Cosmotron and the Bevatron is related to parity non-conservation, i.e., left-right or mirror symmetry violation. Breaking mirror symmetry means that the image would be different from the original, e.g., placing pears in front of a symmetry-broken mirror might result in an image showing apples. The story started in 1949, Rosemary Brown of the Powell group [20, 21], identified on a emulsion plate, exposed to cosmic rays, an unexpected decay event. A hitherto unknown object, given the name k particle, with a mass of about $900~{}m_{e}$ decayed by two distinct theta and tau modes. For a long time, the validity of those data and a few rare follow-up observations by other groups were considered skeptically. By 1953, in total, 11 k particle events were known. When the Cosmotron and the Bevatron delivered their respective first beams, campaigns were started on that issue, rapidly increasing the number of k particle observations and the related theta- and tau-decays, now obtained under controllable laboratory conditions. The two separate decay modes were identified as given by two and three pions, respectively [22]. When Chinowsky and Steinberger [^3] Jack Steinberger (1921-2020) was a recipient of the 1988 Nobel Prize in Physics, along with Leon M. Lederman (1922-2018) and Melvin Schwartz (1932-2006), for the discovery of the muon neutrino) in 1954 could prove experimentally the intrinsic negative parity of pions [23], it was clear that the k particle suffers a parity-violating decay. Shortly after, the famous Wu experiment, performed by Chien-Shiung Wu (1912-1997), cleared skies by showing convincingly that weak interaction is indeed violating parity conservation [24] [^4] To the surprise of many Chien-Shiung Wu did NOT receive the Nobel prize for her pioneering Experimentum Crucis, thereby sharing the fate of many other women in science.. At that time, nomenclature was changed to $K$ meson or kaon. A posteriori, it’s evident that the puzzling $k$ particle decay was indeed the first observation of the weak decay of the kaon. It is noteworthy to mention that in 1954 Chinowsky and Steinberger also showed that charged and neutral pions have slightly different rest masses [25].
Since their discovery, neutral kaons $K^{0}$ together with their anti-particle counterparts $\bar{K}^{0}$ have been a laboratory for research on fundamental aspects of particle physics. They were the first hadrons showing oscillation phenomena by the formation of short-lived ($K_{S}$, $t_{1/2}\approx 8\cdot 10^{-11}~{}\mathrm{s}$) and long-lived ($K_{L}$, $t_{1/2}\approx 5\cdot 10^{-8}~{}\mathrm{s}$) superpositions of the flavor (strangeness) eigenstates $K^{0}$, $S=-1$ and $\bar{K}^{0}$, $S=+1$, respectively. The differences in lifetime are due to the differences in the phase space available for the two-pion and the three-pion decay channels: the larger two-pion phase space causes a large decay width and shorter lifetime than the three-pion decay.
[…]
As early as 1967, Andrey Dmitrievich Sakharov (1921-1989) [^5] In later years Sakharov engaged in human rights activities, turning into a Soviet dissident, and received in 1975 the Nobel prize for Peace for emphasizing human rights around the world. published in [31] the famous three Sakharov conditions on the cosmological matter-antimatter asymmetry problem. The second paradigm demands violation of C and CP symmetry. The modern understanding of CP violation on the microscopic level in hadrons is related to a CP-violating phase factor in the Cabibbo–Kobayashi–Maskawa (CKM) matrix [32] (named after Nicola Cabibbo (1935-2010), Makoto Kobayashi (born 1944), and Toshihide Maskawa (1940-2021))). Kobayashi and Maskawa were awarded half of the 2008 Nobel Prize in Physics for the discovery of the origin of the broken symmetry which predicts the existence of at least three families of quarks in nature. The CKM matrix describes the mixing of quarks under weak interactions, which is also relevant in another open nuclear physics problem, namely neutrinoless double beta decay. Taking together, these research areas are connecting sub-nuclear and cosmological phenomena. The present status about conserved symmetry is that time reversal combined with CP symmetry is conserved, forming together the triple-symmetry CPT.
As a side remark it might be of interest that Richard Feynman (1918-1988) is said to have placed at a conference in 1955 a 50 US Dollar bet against parity violation [33], underlining how unreal such a conjecture appeared at the time of the first rumors of an experimental signature. Feynman is eminently known for his contributions to the development of quantum electrodynamics for which he received the Nobel Prize in Physics in 1965 jointly with Julian Schwinger (1918-1994) and Shinichiro Tomonaga (1906-1979). Moreover, Feynman’s name is engraved forever in the Feynman diagrams which are the indispensable tools for modern QFT research in any theoretical context.
Figure 4: Photograph of the cloud chamber event recording the decay of a heavy neutral particle of two-pronged fork V-particle pattern. The cloud chamber of size $(240\times 400\times 460)~{}\mathrm{mm^{3}}$ is on exhibition in the science museum of the Imperial College of London, Great Britain. Photograph taken from Ref. [15].
2 Charm and Beyond
The focus of this article is on the physics of hadrons made of up ($u$), down ($d$), and strange ($s$) valence quarks and anti-quarks, respectively. Charm physics is discussed in due detail elsewhere in this Encyclopedia [34]. However, since the charm quark is the partner of the strange quark in forming together the second generation of the quark family, we briefly give an overview of the exciting history of charm discovery.
As early as in 1964, the idea came up that there might be more flavors than $\{u,d,s\}$. James Bjorken (1934-2024) and Sheldon Glashow (born 1932) hypothesized the existence of another quark, which they named charm ($c$). As Glashow declared in 1976 in a newspaper article in New York Times the charm quark received its name because of the symmetry it brought to the Subnuclear world as cited in [35]. Experimentally, a signature of charm was observed first in 1974 in two independent experiments at BNL by Samuel C. C. Ting (born 1936) and collaborators [36] and at SLAC by the group of Burton Richter (1931-2018) [37]. Both teams found at an energy of $3.1~{}\mathrm{GeV}$ a narrow resonance with a width of only $5~{}\mathrm{MeV}$. A particular strong confirmation for the existence of charm was that the BNL group used the purely hadronic nuclear reaction $pBe\to e^{+}e^{-}+X$ while the SLAC group observed the new resonance in purely leptonic $e^{+}e^{-}$ annihilation reactions. Thus proving the ubiquity of the phenomenon. The BNL group gave that state the name $J$, the competitors from SLAC chose $\psi$, and finally, as a compromise and honoring the work of both teams, the two labels emerged into the official name $J/\psi$.
Ting and Richter met on November 15, 1974; they issued a press release about their discoveries. The charm story then took up speed: On November 21, 1974, the SPEAR@SLAC collaboration already reported the observation of $\psi^{\prime}(3700)$, an excited state of the $J/\psi(3100)$. Both particles are charmonium states, composed of $c\bar{c}$ valence quarks. The modern spectroscopic notation is $J/\psi(1S)$ and $\psi^{\prime}(2S)$, following the constituent quark model and indicating that $\psi^{\prime}$ is understood as a radial (compressional) excitation of $J/\psi$. The latest mass values are $M_{J/\psi}=3096.900\pm 0.006~{}\mathrm{MeV}$ and $M_{\psi^{\prime}}=3686.097\pm 0.010~{}\mathrm{MeV}$ [38].
The two group leaders Ting and Richter were awarded the 1976 Nobel prize, honoring the great achievement of the two teams for for discovering the subatomic $J/\psi$ particle. Their epochal discoveries became known as the November Revolution of particle physics.
In 1986, Matsui and Satz predicted $\ldots$ that $J/\psi$ suppression in nuclear collisions should provide an unambiguous signature of quark-gluon plasma formation [39]. Since then, the $J/\psi$ plays a special role in the QGP searches in heavy-ion collisions at ultra-relativistic energies at RHIC and the LHC.
A multitude of hidden charm charmonium states and open charm mesons and baryons have been observed, containing a charmed quark or anti-quark to which $\{u,d,s\}$ partners and the respective anti-quarks are attached. There is strong evidence that charmed hadrons may develop exotic configurations, e.g., tetra-quark states or molecular-like states formed by D mesons. These so-called $X,Y,Z$ states fall out of the systematics of the otherwise very successful quark models for quarkonia. Their spectroscopy is discussed elsewhere in this Encyclopedia [40].
Corresponding hadron configurations are found for bottom ($b$) quarks and have to be expected in the top ($t$) quark sector. Bottom and top quarks together represent the third generation of the quark family. In Fig. 5, a table summarizing the three quark generations and their properties is shown.
3 Theoretical Approaches and Interpretation of Data
1 Aspects of QCD
QCD is the fundamental theory of strong interactions and thus of hadrons666 Lev Borisovich Okun (1929-2015) introduced the term hadron (ancient Greek for heavy or strong) in a plenary talk at the 1962 International Conference on High Energy Physics.. The beauty and peculiarities of QCD as a non-Abelian quantum gauge field theory of intrinsically non-perturbative character will be discussed elsewhere in this Encyclopedia. Here, an overview is given on the aspects that are of special interest and relevance for hadron physics and the production of hadrons.
In the early days of nuclear and elementary particle physics, the plethora of particles observed in cosmic-ray and high energy experiments made improbable their being elementary particles. Obviously, a new approach was necessary, allowing for bringing order into the data. In 1961, Yuval Ne’eman (1925-2006) introduced the classification of hadrons through the SU(3) flavor symmetry, exploiting the representation of the SU(3) group in terms of eight linearly independent 3-by-3 matrices and three elementary degrees of freedom, later named the Eightfold Way [41, 42]. Similar group-theoretical descriptions were invented independently by Murray Gell-Mann (1929-2019), who finally formulated QCD [43]. Murray Gell-Mann was awarded the Nobel Prize in Physics in 1969 for his contributions and discoveries concerning the classification of elementary particles and their interactions.
In parallel to and independently of Ne’eman and Gell-Mann, George Zweig (born 1937) developed during his stay as a post-doc at CERN his own version of a SU(3) scheme of strongly interacting particles [44, 45]. His contributions to QCD are honored in the OZI rule, accounting, e.g., for the hindered decay of the $\varphi(1020)$ vector meson. Thus, by 1964, two competing group theoretical approaches were available, relating the abstract mathematical structure of non-commuting Lie algebra (named after the mathematician Sophus Lie (1842-1899)) to elementary particle physics. Although at that time the authors insisted on essential differences between their theories, there were obvious overlapping similarities. Both approaches were in need of introducing three elementary entities in order to place mesons and baryons on the same footing. While Zweig chose the name aces, inspired by the joker cards of card games, Gell-Mann opted for quarks [^7] Three quarks for Muster Mark” is a line from James Joyce’s novel “Finnegans Wake.” (It took for James Augustine Aloysius Joyce (1982-1941) 16 years to complete this experimental novel.) The phrase is notable because it provided the name “quark” for a fundamental subatomic particle, chosen by Murray Gell-Mann. Gell-Mann initially considered the word “kwork” before encountering the Joyce passage and finding it more fitting. While the original line seems to rhyme “quark” with “Mark,” Gell-Mann proposed a justification for pronouncing it ”kwork” by associating it with the pub owner’s call of “Three quarts for Mister Mark. - which finally was adopted by the community. In Figure (Table) 5, the quarks and their properties are listed. The discovery of new particle species ($J/\psi$) in 1974 enforced the introduction of the charm (c) quark and later discoveries of particles required to extend the list of QCD flavors to bottom (b) and top (t) quarks.
Figure 5: Properties of quarks of the first $(u,d)$, second $(c,s)$, and third $(t,b)$ generation [38]. As a generalization of isospin, originally introduced for $\{u,d\}$ systems and used above to distinguish proton and neutrons by the eigenvalues of the third component of the isospin operator $I_{3}=\pm\frac{1}{2}$, heavier quarks are characterized by the flavors $c,s,t,b$ with values $0,1$. In this chapter, only production processes involving hadrons built of the first and second quark generation are considered. The figure is adapted from Ref. [46].
The works of Yuval Ne’eman, Murray Gell-Mann, and George Zweig were the long-awaited breakthrough towards a new understanding of hadrons in terms of a few elementary degrees of freedom, given by the massive quark Fermion fields and by gluon gauge vector fields, which are the force carriers of strong interactions. Different from the photons of QED, gluons carry color charges and interact with each other.
An immediate consequence of the group-theoretical approach is to introduce order schemes in terms of multiplets which are defined by the invariants of the respective mathematical group. In Fig. 6, the lowest octet and decuplet SU(3) multiplets of baryons, composed of up (u), down (d), and strange (s) quarks, are shown by being arranged in a diagram constructed by the two group invariants hypercharge, $Y=B+S$, defined by the baryon number $B$ and strangeness $S$ and the third component of the SU(2) isospin operator, $I_{3}=Q-\frac{1}{2}Y$, where $Q$ is the baryon charge in units of the elementary charge $e$. The algebra underlying multiplet diagrams as in Fig. 6 was described by de Swart [47] (Johan de Swart (1931-2014)) in a comprehensible, compact manner and has been used widely until today.
Although group theory does not provide by itself absolute mass scales, the mathematical structures predict highly valuable sum rules relating the masses within multiplets. In [48], Murray Gell-Mann exploited those relations to identify the already known two-pion rho-meson ($\rho$) resonance to be part of a vector meson octet. He also predicted that there must be an isoscalar partner of the rho-meson decaying into three pions, which he named omega-meson ($\omega$). Moreover, at the International Conference on High-Energy Nuclear Physics, Geneva, 1962, Gell-Mann used the mass relations to predict that the still missing last decuplet member of strangeness $S=-3$, i.e., the $\Omega^{-}$ hyperon, would have the mass $M^{(theo)}_{\Omega}=1676~{}\mathrm{MeV}$. Two years later, the particle was indeed found at Brookhaven by Barnes et al. [49, 50] at a mass of $M^{(exp)}_{\Omega}=1674\pm 3~{}\mathrm{MeV}$ [^8] The BNL 80-in. hydrogen bubble chamber was exposed to a mass-separated beam of $5.0~{}\mathrm{BeV/c}$ K mesons at the Brookhaven AGS. About 100,000 pictures were taken containing a total K track. The analysis shown a single event associated with the $\Omega^{-}\to nK^{0}\pi^{-}$ decay.. That extraordinarily good agreement gave strong support and confidence to the group-theoretical SU(3) approach.
At the same time as Gell-Mann but independently, Susumo Okubo (1930-2015) was pursuing at Rochester his own studies of unitary symmetry in strong interactions, leading to mass relations of the same type [41, 51, 52]. Since then, the SU(3) mass relations are known as Gell-Mann-Okubo (GMO) mass formulas [^9] Deviations from the GMO mass formula predictions are often important hints for mass shifts form dynamical mixing..
One of the central predictions of QCD is the parton structure of hadrons. Once that conjecture was confirmed by experiment in the early 1970’s [53], QCD evolved into the nowadays accepted standard model of strong interaction physics. QCD describes hadrons as systems of quarks, electrically charged spin-$\frac{1}{2}$ fermions with flavor, color and isospin, and gluons, colored spin-$1$, self-interacting vector gauge fields. Outstanding features of QCD are quark confinement and color-neutrality of physical particles, running coupling constants, and asymptotic freedom. Renormalization group techniques and regularization methods are indispensable parts of the QCD mathematics toolbox. Spontaneous symmetry breaking is a defining mechanism of QCD. The complexities of the non-Abelian structures of QCD (named after Niels Henrik Abel (1802-1829)), leading to higher order non-linear self-interactions, inhibit solving the deeply nested field equations in closed form. Only in the ultra-high energy regime of asymptotic freedom, where perturbative treatments become possible, leading order results can be derived analytically. The electro-weak theory of Salam, Weinberg, and Glashow, and QCD are the constituencies of the Standard Model (SM) of elementary particle physics. Mohammad Abdus Salam (1926-1996) shared the 1979 Nobel Prize in Physics with Steven Weinberg (1933-2021) and Sheldon Glashow (born 1932) for his contribution to the electroweak unification theory.
For a long time, the missing link was a mechanism which could explain within the SM the origin of mass. In 1964, Peter Higgs (1929-2024) proposed a mechanism for mass generation which was compatible with the SM but postulated the existence of a boson (named after Satyendra Nath Bose (1894-1974)) multiplet [54, 55]. In 2012, finally the (neutral) Higgs boson $H^{0}$ (colloquially called “God Particle”) was observed at about the expected mass $m_{H}=(125.20\pm 0.11)~{}\mathrm{GeV}$ [38] at the LHC at CERN. Since then, the SM is a closed theory, although lacking a link to gravitation. For this work, Higgs received the Nobel Prize in Physics in 2013, which he shared with François Englert (born 1932).
Figure 6: The first two baryon SU(3) flavor multiplets are given by an octet (left) and a decuplet (right). The valence quark content of the baryons is indicated explicitly. The vertical axes are representing the hypercharge $Y=S+B$, given by the strangeness $S$ and baryon number $B$; the horizontal axes indicate the third component $I_{3}=Q-\frac{1}{2}Y$ of the isospin $I$ includes also the charge number $Q$. The group theoretical background and construction of these kind of diagrams is discussed in depth in textbooks, see, e.g.[56].
Mesons and baryons may be considered as the asymptotic mass eigenstates of QCD in the low energy limit. However, they are of fundamentally different character. Considered as systems of quarks and gluons, mesons are formed by an even number of valence quarks, while baryons are defined as the particles containing an odd number of valence quarks. In both cases, an arbitrary amount of gluons is allowed in addition. Since quarks are Fermions of spin $s=\frac{1}{2}$ and positive parity, baryons carry half-integer spin quantum numbers $J_{B}=\frac{1}{2},\frac{3}{2},\frac{5}{2},\ldots$, while mesons have integer spin quantum numbers $J_{M}=0,1,2\ldots$. Moreover, baryons play a special role in cosmology because they account as protons and neutrons together with a minor contribution of electrons for the matter content of the Universe.
Figure 7: The lowest meson SU(3) flavor nonets of pseudo-scalar (left), vector (center), and scalar mesons (right) as functions of hypercharge Y and isospin I3. Different to the baryons, mesonic two-quark systems of $\{u,d,s\}\otimes\{\bar{u},\bar{d},\bar{s}\}$ structure arrange into SU(3) octets and accompanying singlets. The isoscalar (charge neutral) octet mesons obtain an isoscalar singlet partner. The singlet mesons are included into the octet diagrams, giving rise to the octet-singlet pairs $\{\eta,\eta^{\prime}\}$, $\{\omega,\phi\}$, and $\{\sigma\simeq f_{0}(500),\sigma^{\prime}\simeq f_{0}(980)\}$ in the pseudo-scalar, vector, and scalar meson sectors, respectively. Singlet and octets are mixing by, in principle, arbitrary mixing angles. A widely used choice is Ideal Mixing, chosen such that the isoscalar octet mesons $\eta,\omega,\sigma$ do not contain a $s\bar{s}$ component. Electromagnetic effects lead to $\omega-\rho$ mixing and correspondingly $\sigma-\delta^{0}$ mixing has to be expected. Scalar mesons may contain multi-pion. i.e., multi-$q\bar{q}$ components. The complexities of scalar mesons are especially visible in the isovector scalar meson $\delta\simeq a_{0}(980)$. $a_{0}(980)$ is decaying preferentially into $\eta\pi$, finally ending in three pions, but also $K\bar{K}$ and $\eta^{\prime}\pi$ decays are reported [38]. The physical counterpart of the $\kappa$ meson is the $K^{\ast}_{0}(700)$ meson of a notoriously uncertain pole structure [38]. Thus, the physically observed mesons are not identical with the bare SU(3) mesons.
The seemingly simple picture of hadrons emerging at first sight from QCD is in fact deceiving because the complex field-theoretical dynamics lead in fact to a large variety of spectroscopic configurations. The basic constituents, quarks and gluons, are interacting in a manifold manner which will lead to a variety of intrinsic configurations forming the inner confined core of a hadron. A selection of expected elementary configurations is displayed in Fig. 8. The core region of a physical hadron will contain a mixture of those configurations, i.e., hadron wave functions are typically quantum mechanical superpositions of multi-quark, quark-gluon, gluon-gluon, etc. components - all embedded on top of a highly dynamical vacuum. The core region, albeit confined to an incredibly small volume of about $(0.5\times 10^{-45})~{}\mathrm{m^{3}}$, is only part of the complete hadron. The quark-gluon core polarizes the vacuum and, as a consequence, is surrounded by a cloud of virtual particles, mainly consisting of (pairs of) pions and other mesons.
Asymptotic freedom on the one side and quark confinement on the other side are defining parts of strong interaction physics. Gluon self-interactions and effects from the spontaneously broken chiral symmetry are important ingredients of constituent quark and hadron masses. The highly non-linear nature of QCD inhibits analytical or perturbative solutions for most parts of the accessible energy region, except for the highest energies as, e.g., reachable at the LHC at CERN. The mathematical complexities of QCD at energies relevant for hadron structure, spectroscopy, and interactions of hadrons can be treated only numerically. LQCD is the method of choice, being used successfully to explore mesons and baryons by QCD principles.
Figure 8: Illustration of elementary quark-gluon configurations of mesons (left) and baryons (right). The object in the center is a glueball, i.e., an extremely exotic hadron consisting solely of self-interacting gluons.
4 Hadron Spectroscopy
1 QCD and Phenomenology
While the high energy sector of QCD is well understood on the basis of perturbative QCD, the low-energy limit is still a field with many open questions, reflecting the highly nontrivial task of handling a non-perturbative quantum field theory. Formally, the QCD path integral is the proper starting point. As discussed elsewhere in this encyclopedia, two alternative numerical schemes are currently pursued: In LQCD, a set of coupled correlation functions is derived from the path integral and propagated on a Euclidean (named after Euclid (about 325 — 265 BC), an ancient Greek mathematician) space-time grid where masses are determined by the numerically produced pole structures [57]. One of the first successful LQCD descriptions of the mass spectrum of the octet mesons and the octet and decuplet baryons, respectively, was achieved by the Budapest-Wuppertal Collaboration [58, 59]. The mass spectrum is displayed in Fig. 9. Functional methods derive sets of coupled Dyson-Schwinger equations (DSE) (named after Freeman John Dyson (1923-2020) and Julian Schwinger) by performing variational functional derivatives of the path integral [60].
While LQCD and DSE approach hadron spectroscopy from the inner sector of the quark core of hadrons, effective field theories and phenomenological reaction models put their focus on the properties reflecting the nature of hadrons as dynamically generated composite states. Compositeness is another aspect of the dualism hidden in baryons and mesons as visible in the observed resonances in scattering and production experiments. Until today, phenomenological approaches based on an underlying covariant field theory are an important source of spectroscopic information on the quantum numbers of resonances, their formation, and their decay. The models are apt for large-scale numerical coupled channels calculations utilizing the partial wave formalism. Meson production on the nucleon is described through the excitation of baryon resonances in photo- and meson-induced reactions on the nucleon. Baryon resonances are identified through their traces left as complex energy poles in the partial wave scattering matrix elements, which in many cases are strongly affected by coupled channel effects.
At masses above the pion-nucleon threshold, baryon resonances have to be considered as superpositions of molecular-like, loosely bound, or unbound meson-nucleon configurations. They are in competition with the genuine QCD-type confined quark core configurations, illustrated in Fig. 8. In addition, the QCD core is shielded from the exterior by a virtual $q\bar{q}$ cloud of transient mesons. Very likely, also glue-ball-like configurations will contribute as indicated by the large differences between the current and the constituent quark masses. A given resonance will be composed of an arbitrary mixture of two building blocks. In some cases, the quark-type configurations will dominate; in others, the molecular configuration may prevail. An intensively and controversially debated candidate for the latter type of states is the $\Lambda^{\ast}(1405)$ ($S=-1$) resonance, which falls out of the systematics of quark models but is characterized by a pronounced pole structure in the complex plane. The chiral unitary model [61] predicted a two-pole structure from a superposition of $\pi\Sigma$ and $\bar{K}N$ components. Years later, improved experimental and theoretical studies concluded that $\Lambda^{\ast}(1405)$ is dominantly a $\bar{K}N$ composite; however, without fully excluding $\pi\Sigma$ admixtures [62].
5 The Constituent Quark Model and Hadron Spectroscopy
The discovery of resonance phenomena in the early pion-nucleon scattering experiments cast strong doubts on the initial assumption that nucleons were structure-less elementary point particles, as was assumed - and since then confirmed - for the electron.
Table 1: The mass spectrum of constituent quarks. the masses are derived from the data od meson and baryon masses such that the experimental masses are reproduced by the sum of constituent quark masses. The quark generations are indicated by separating lines. For further details see Ref. [63].
Constituent Quark Mass [MeV]
Up (u) 336
Down (d) 340
Strange (s) 486
Charm (c) 1,550
Bottom (b) 4,730
Top (t) 177,000
Once QCD had established quarks and gluons as the constituents of hadrons, the interest in nucleon spectroscopy was inevitable. The CQM, going back to Isgur and Karl [64, 65, 66, 67, 68, 69, 70, 71], was an important step forward to understand and predict hadron spectra on phenomenological grounds. Mesons are described by quark-antiquark ($q\bar{q}$) configurations and baryons as states of three valence quarks ($qqq$). The gluon component in hadron masses - actually accounting for the major part of the masses - is taken into account by assigning to the quarks effective constituent masses which are derived empirically from data. The constituent quarks interact by a Coulomb-like static potential [^10] Charles-Augustin de Coulomb (1736-1806) derived first the $1/r^{2}$ behavior of forces between electric charges, corresponding to a potential $\sim 1/r$., modeling gluon exchange, and are enclosed by a confinement potential, typically chosen as a power law in the radial coordinate. Spin- and flavor-dependent residual interactions are used to describe the fine structure of hadrons. Modern versions of the Cornell potential, originally introduced in the early 1970s by Eichten et al. [72] are a standard tool for the modeling of hadrons, firmly established in spectroscopic studies of heavy mesons in the charm and bottom sectors. The CQM is discussed in breadth elsewhere in this Encyclopedia [63], including also a concise derivation on QCD grounds.
Figure 9: The light hadron spectrum of QCD as obtained by QCD. Horizontal lines and bands are the experimental values with their decay widths. QCD results are shown by solid circles. Vertical error bars represent combined statistical systematic error estimates. The pion ($\pi$), $S=-1$ kaon ($K$), and the $S=-2$ cascade baryon $\Xi$ have no error bars, because they were used to set the $u,d,s$ quark masses and the overall scale, respectively. The figure is adapted from Ref. [58].
The CQM plays an important role in counting, ordering and classifying hadron states. For example, the CQM wave functions of baryons composed of three fermions are given by
$|\Phi_{3q}\rangle=|\psi_{color}\rangle\otimes|\psi_{flavor}\rangle|\psi_{spin}\rangle|\psi_{space}\rangle\>,$ (1)
which in total must be antisymmetric. Color neutrality demands that the color-component is a singlet state which is antisymmetric by definition and construction. Thus, the remaining flavor-spin-space part must be symmetric which implies that the spin-flavor and the space part belong to the same symmetry class, i.e., both are symmetric or antisymmetric or of mixed symmetry. As explained in detail by Giannini [73] and in many textbooks, the symmetry condition imposes strict constraints on the allowed baryon configuration. In the $\{u,d,s\}$ sector, the spin and flavor degrees of freedom can be combined conveniently to form a SU(6) group. Combinatorics leads immediately to a total number of 216 allowed spin-flavor configurations. The respective irreducible representations are arranged into four SU(6) multiplets, given by a 20-plet of antisymmetric states, two 70-plets of mixed symmetry, and a 56-plet of symmetric states. The 20-plet and the 70-plets include flavor singlet states which typically are neglected because of lacking empirical evidence. The remaining irreducible components are flavor octets and decuplets, both of total spin $S=\frac{1}{2}$ and $S=\frac{3}{2}$, respectively. Evidently, a plethora of baryon states will be obtained by combining the SU(6) spin-flavor group structures and space parts provided by the O(3) group.
1 Counting Resonances
Soon it was realized that there was an obvious discrepancy between the number of resonances predicted by theory, as the rather successful constituent quark model, and those identified experimentally. Nathan Isgur (1947-2001) was probably the first who introduced the term missed resonances [74], thus honoring the mismatch between the more than 400 baryon states predicted group-theoretically by $SU(6)\times O(3)$ multiplet rulings (three 70-plets and four 56-plets) - about the same number is obtained by LQCD - and the number of experimentally verified resonances plus proton and neutron which amounts to a little more than 100 safely confirmed states PDG2024 [38]. The problem of missing resonances is a major issue of baryon spectroscopy. Final answers about the number of excited states of the nucleon and their spectral properties are still pending. Solutions are sought for both experimentally and theoretically.
The current understanding of the internal quark-gluon composition of a hadron is illustrated in Fig. 8. The observed states, however, are the result of additional interactions with the set of meson-baryon scattering states of the same total quantum numbers. Hence, the complete picture of a hadron necessarily will lead to a multi-configuration problem, superimposing the internal, confined quark-gluon components and the external meson-baryon configurations, either as a virtual polarization cloud or as on-shell continuum channels. Obviously, quark models by themselves will account for such additional dynamical effects only in a weak coupling limit where mass shifts and other effects can be subsumed into effective phenomenologically determined parameters.
In general, polarization effects, discussed below, induce self-energies which shift the bare QCD configurations in mass, and redistribute the spectroscopic strength i over the eigenstates of the coupled system. Formally, the eigenenergies of the interacting systems will be found to be moved into the complex plane where the imaginary parts define the decay and the lifetime of the polarized states. Thus, baryons heavier than the nucleon-pion system acquire rather short lifetimes because they may decay either by strong (and, to a minor degree, electromagnetic) interactions or, as in the case of hyperons, by weak interactions.
The spin-isospin multiplet of $\Delta_{33}(1232)$ states, belonging to the $J^{P}=\frac{3}{2}^{+}$-baryon decuplet, is a typical case: As a QCD state, the Delta-resonance is understood as a spin-isospin vector ($\Delta S=1=\Delta I$) excitation of the nucleon by reorienting the spin and isospin of at least one quark, resulting in the $(J,I)=(\frac{3}{2}\frac{3}{2})$ configuration. That particular configuration is coupled, however, strongly to the continuum of pion-nucleon $P$-wave scattering states, which induces a rapid decay within a time span of $t_{1/2}\sim 10^{-23}~{}\mathrm{sec}$, corresponding to a spectral distribution of full width at half maximum (FWHM) $\Gamma_{\Delta}\simeq 120~{}\mathrm{MeV}$. In contrast to other baryon resonances, the Delta-resonance is prominently excited in practically all types of reactions, from virtual and real photo-excitation by electrons and photons to neutral and charged current neutrino-nucleon and neutrino-nucleus reactions, as well as in hadron scattering both on elementary proton targets and on nuclear targets. The description of those multi-channel and multi-configuration phenomena is the domain of the coupled channels methods discussed in the following sections.
2 Notations
The CQM spectroscopic notation is $|B{}^{2S+1}L_{J}\rangle_{\sigma}$ where $B=N,~{}\Delta$ denotes an octet or a decuplet, $L=S,~{}P,~{}D\ldots$ denotes the total orbital momentum, $|L-S|\leq J\leq L+S$ is the total angular momentum. The flavor symmetry character is specified by $\sigma=A,M,S$ for Asymmetric (A), Mixed (M), and Symmetric (S) .
Traditionally, a different notation is used in meson-nucleon spectroscopy where states are classified by $L_{2I,2J}$, according to their pion-nucleon partial wave quantum number $L=S,~{}P,~{}D\ldots$, twice the total pion-nucleon isospin $I=\frac{1}{2},\frac{3}{2}$ and twice the total spin $J$. In that notation $I=\frac{1}{2}$, nucleon-like states are denoted by $P_{11}$, $P_{13}\ldots P_{12J}$, e.g., the nucleon is $P_{11}(940)$ and $P_{11}(1440)$, which is the Roper resonance (discovered in 1964 by L. David Roper (born 1935)). Thus, the resonance (pole) mass is added for distinguishing the resonances within the same partial wave.
The non-strange $I=\frac{3}{2}$ decuplet states are labeled by $\Delta_{32J}$, where the best-known case is the lowest Delta-isobar resonance, $\Delta_{33}(1232)$ which was discovered by Enrico Fermi and his group in 1952. For pion-nucleon S-wave configurations and the higher partial waves $L\geq 2$, the standard spectroscopic notation $L_{2I,2J}$ is used, combined with the resonance mass. There are ambiguities in the choice of resonance masses. From a theoretical point of view, unstable, decaying states are characterized by the spectroscopic pole in the complex plane, including the full information on self-energies from the interactions with the surrounding continuum. However, frequently, the so-called Breit-Wigner mass (named after Gregory Breit (1899-1981) and Eugene Paul Wigner (1902-1995) [^11] Wigner received the Nobel Prize in Physics in 1963 for his contributions to the theory of the atomic nucleus and the elementary particles, particularly through the discovery and application of fundamental symmetry principles. ) is used instead [75].
The international Particle Data Group (PDG), founded in 1957 by Murray Gell-Mann (1929-2019) and Arthur Hinton Rosenfeld (1926-2017) (the latter named “Godfather of Energy Efficiency”) [76]) uses a slightly different notation, as encountered in the latest PDG edition [38]: The proton and the neutron are distinguished by $p$, $I_{3}=+\frac{1}{2}$ and $n$, $I_{3}=-\frac{1}{2}$, thus accounting for the slight breaking of isospin symmetry in the nucleon iso- doublet as reflected in the p-n mass difference of about $1~{}\mathrm{MeV}$. Higher mass members of the $I=\frac{1}{2}$ nucleon family are denoted by $N(M_{n})J^{\pi}$, where $\pi=(-1)^{(}L+1)=\pm(1)$ is the parity of the state, composed of the orbital angular momentum $L$ of the partial wave and the intrinsic (negative) parity of the pion. $M_{n}$ is the resonance mass, also specifying the position of the state in the spectral sequence.
The strangeness sector, $S=-1$ and $S=-3$ iso-singlet ($I=0$) baryons are denoted by $\Lambda(M_{k})J^{\pi}$ and $\Omega(M_{k})J^{\pi}$, respectively. Correspondingly, the iso-triplet ($I=1$) baryons of strangeness $S=-1$ and the iso-doublet ($I=\frac{1}{2}$) of strangeness $S=-2$ are classified by $\Sigma(M_{k})J^{\pi}$ and $\Xi(M_{k})J^{\pi}$, respectively. By historical reasons, the generic names Lambda ($\Lambda$), Sigma ($\Sigma^{0,\pm}$), Cascade ($\Xi^{0,-}$), and Omega ($\Omega^{-}$) are in use for the lowest octet and decuplet hyperons.
By obvious reasons, most of the measurements are performed on proton (hydrogen) targets. Neutron data are obtained mainly on deuterium targets, used as a surrogate for non- existing elementary neutron targets and taking advantage of the loosely bound deuteron with a large distance between the constituents. The proton as a free particle was identified first by Ernest Rutherford (1871-1937) in 1920, but already in 1908 he was awarded the Nobel Prize in Chemistry for his investigations into the disintegration of the elements, and the chemistry of radioactive substances. The neutron was discovered by James Chadwick (1891-1974) in 1932 who received the Nobel Prize in Physics in 1935 for his discovery of the neutron.
6 Coupled Channels Methods to Meson Production and Hadron Spectroscopy
1 Interacting Hadrons
The quark-gluon configurations displayed in Fig. 8 are building blocks of the QCD core and could be taken as a basis for extended studies. The core eigenstates are obtained by diagonalization within the quark-gluon configuration space, which is demanding but not prohibitive. The result is a spectrum of energy-sharp, discrete eigenstates because of confinement. CQM and bag models of hadron structure are coming closest to that kind of picture. However, for describing physical hadrons realistically, it has to be taken into account that these model states are embedded in the continuum of hadronic scattering states [^12] In molecular, atomic, and nuclear physics such states are known as bound states in the continuum (BIC)..
As theoretical constructs, the bare quark-gluon states correspond to closed channels with inseparable constituents under normal environmental conditions. The (multi-)hadron-hadron scattering states are open channels with asymptotically free particles. Depending on the case, they are composed either of pure meson and baryon-antibaryon for $B=0$ systems, or of mixed meson-baryon scattering states in $B\neq 0$ objects. If there would be 6-quark objects, e.g., the notorious and experimentally still searched for $S=-2$ H-dibaryon hypothesized by Jaffe [77], baryon-baryon channels would be involved as well. The observation of the non-strange $d^{\ast}(2380)$ dibaryon, probably a double-$\Delta^{0}$ configuration, was reported by the WASA-COSY collaboration [78]. Theoretical studies find rather compact d-baryon wave functions [79]. In either case, the set of discrete channel quantum numbers, defined by the coupled irreducible representations of the hadronic constituents, must coincide with those of the core [^13] It’s worth noting that the best known and stable dibaryon is the deuteron, although not of the compact stature as expected for exotic states.. Functional methods and LQCD will in principle account for such effects, as the good reproduction of the experimental mass spectrum in Fig. 9 shows.
The basic mechanisms governing the interplay of quark-gluon and hadron-hadron components of hadron wave functions are easily understood in Hamiltonian formulation (named after William Rowan Hamilton (1805-1865)). The description of physical hadrons demands solving a coupled channels problem.
2 Interactions of QCD Core and Hadron Scattering Configurations
Consider a QCD core state $|Q\rangle$ of bare mass $M_{Q}$ which is coupled by the interactions $V_{QC}=V^{\dagger}_{CQ}$ to a set of open hadron scattering channels $C$ of total channel mass $M_{C}=\sum_{h\in C}m_{h}$, given by the sum over the masses $m_{h}$ of all asymptotically free hadrons $h=p,n,\pi\ldots$. Standard projection methods allow us to recast the set of coupled equations by a fully equivalent dispersive, non-Hermitian polarization self-energy operator [^14] An operator $A$ is Hermitian if $A=A^{\dagger}=(A^{T})^{*}=(A^{*})^{T}$ does not change under transposition ($T$) and complex conjugation (*). A real number $x$ is Hermitian, a complex number $z=x+iy$ in non-Hermitian. Any non-Hermitian operator can be decomposed in a Hermitian and an anti-Hermitian operator (named after Charles Hermite (1822-1901)). acting on the quark-gluon core:
$\Sigma_{Q}(w_{\alpha})=\sum_{C}\int\frac{d^{3}q}{(2\pi)^{3}}\frac{|V_{QC}(\mathbf{q})|^{2}}{w_{\alpha}-w_{C}(q)+i\eta}\>,$ (2)
where $\eta\to 0+$ and $w_{\alpha}=\sqrt{s_{\alpha}}$ denote the invariant energy brought in by an external probe $\alpha$.
By means of the Cauchy formula (named after Augustin-Louis Cauchy (1789-1857)), the propagator is decomposed into a principal value real part and an imaginary part given by a Dirac $\delta$-distribution. Following a widely used notation, the self-energy operator is written as
$\Sigma_{Q}(w_{\alpha})=\Delta_{Q}(w_{\alpha})-i\frac{1}{2}\Gamma_{Q}(w_{\alpha})\>.$ (3)
The real part is
$\Delta_{Q}(w_{\alpha})=\sum_{C}\int\frac{d^{3}q}{(2\pi)^{3}}c|V_{QC}(\mathbf{q})|^{2}\frac{\mathcal{P}}{w_{\alpha}-w_{C}(q)}\>$ (4)
to be evaluated as a Cauchy Principal Value $\mathcal{P}$ (named after Augustin-Louis Cauchy). The imaginary part, expressed by the width
$\Gamma_{Q}(w_{\alpha})=2\sum_{C}\Theta(w_{\alpha}-M_{C})\rho_{C}(q_{\alpha C})\langle|V_{CQ}(q_{\alpha C})|^{2}\rangle\>$ (5)
is determined by the sum of the residues at the poles of the propagator appearing at the equivalent on-shell energy of the intermediate channels. For two-body channels, $M_{C}=M+m$, the invariant on-shell three-momenta are defined by the positive root of
$q^{2}_{\alpha C}=\frac{1}{4w^{2}_{\alpha}}(w^{2}_{\alpha}-(M+m)^{2})(w^{2}_{\alpha}-(M-m)^{2})\geq 0\>,$ (6)
where, different from the real part, only channel states with $M+m<w_{\alpha}$ contribute. The momentum integration has led to the kinematical (phase space) factor
$\rho_{C}(q_{\alpha C})=\frac{q_{\alpha C}E_{M}(q_{\alpha C})E_{m}(q_{\alpha C})}{\pi w_{\alpha}}\>,$ (7)
where $E_{m}(q)=\sqrt{q^{2}+m^{2}}$. The integration over the orientations of the intermediate momentum coupling results in the angle-averaged matrix element
[…]
Including the self-energy, i.e., the coupling to the open hadron channels, the bear (model) states $|Q\rangle$ are transformed to the polarized - in this model the physical - resonances $|R\rangle$. They are determined by a generalized, dispersion eigenvalue problem. The spectrum of resonances is defined by the (complex) roots $\mathcal{M}_{R}$ of the equation
$M_{Q}+\Delta_{Q}(\mathcal{M}_{R})-i\frac{1}{2}\Gamma_{Q}(\mathcal{M}_{R})-\mathcal{M}_{R}=0\>.$ (9)
The solutions are complex eigenvalues $\mathcal{M}_{R}=M_{R}-\frac{i}{2}\Gamma_{R}$, implying a finite lifetime $t_{1/2}=\hbar/\Gamma_{R}$ (using $c=1$) and a mass shift $\Delta_{R}=M_{R}-M_{Q}$.
A closer inspection of the dispersion equations reveals that if the QCD configuration $|Q\rangle$ interacts with $N^{(Q)}_{C}$ channel states $|C\rangle$, Eq. (9) has in total $N_{R}=N^{(Q)}_{C}+1$ solutions of discrete complex eigenvalues $\mathcal{M}_{R_{n}}$. Hence, for each fundamental QCD configuration, the channel coupling will produce a spectrum of $N^{(Q)}_{C}$ satellite states $|R_{n}\rangle$, $n\leq N^{(Q)}_{C}$ plus a solution corresponding to the unperturbed state $|Q\rangle$ of mass eigenvalue $\mathcal{M}_{R_{n}}=M_{R_{n}}-\frac{i}{2}\Gamma_{R_{n}}$. The spectroscopic strength of $|Q\rangle$ is distributed over the spectrum of core-like eigenstates. Typically, but in detail depending on the channel spectrum and the interactions $V_{CQ}$, the state of energy closest to $M_{Q}$ will acquire the largest spectroscopic factor.
In order to maintain analyticity, theoretical approaches should account for the energy dependence of all parts of the self-energies. However, for obvious reasons, little to nothing is known about the intrinsic structures of self-energies because their full knowledge is equivalent to the exact solution of the complete CC problem. In practice, typically the width, i.e., the imaginary part of the channel self-energy, is approximated by functional forms of the proper behavior at the threshold. Close to the threshold of a partial wave with orbital angular momentum $\ell$ and channel momentum $\mathbf{q}$, the coupling matrix elements behave as $V_{CQ}(|\mathbf{q}|\sim q^{\ell}$, implying for the width $\Gamma_{CQ}\sim q^{2\ell}$. Widely used parameterizations are the so-called Blatt-Weisskopf form factors (named after John Markus Blatt (1921-1990) and Victor Frederick “Wiki” Weisskopf (1908-2002)) $F_{\ell}(q/q_{0})$ [80]. These form factors account for the centrifugal barrier effects which determine the behavior of an outgoing scattering wave at the interaction radius. They are given by spherical Hankel functions (named after Hermann Hankel (1839- 1873). In [81], they were generalized to hadron resonance scattering. Explicit formulas and practical applications are discussed, e.g., in Ref. [38].
Once the imaginary part of a self-energy is fixed, the real part is obtained consistently by dispersion theory. The Chew-Mandelstam method (named after Geoffrey Foucar Chew (1924-2019) and Stanley Mandelstam (1928-2016)) accomplishes that goal by evaluating a once-subtraction dispersion integral where the widths are parameterized by Blatt-Weisskopf form factors. In the SAID approach, these effects are incorporated [82].
3 Polarization Self-energies in Hadron Scattering Amplitudes
For understanding the core-channel interplay in hadron production processes, the channel self-energies induced by the core states need to be derived:
$\Sigma_{C}(\omega)=\sum_{Q}V_{CQ}\frac{1}{\omega-M_{Q}}V_{QC}=\sum_{Q}x_{CQ}\frac{\Gamma_{Q}}{\omega-M_{Q}}\>,$ (10)
where $\Gamma_{Q}=\sum_{C}|\rho_{C}\langle V_{CQ}\rangle|^{2}$ and $x_{CQ}=\rho_{C}|\langle V_{CQ}\rangle|^{2}/\Gamma_{Q}$ are the branching ratios, to be evaluated at the appropriate channel momenta. As a consequence, the scattering amplitude of channel $C$ becomes the sum of the scattering amplitudes from the generic channel interactions and the amplitude produced by $\Sigma_{C}$. The latter gives rise to resonances which interfere with the weakly energy-dependent in-channel amplitudes. As a function of energy, the respective partial wave cross sections will show peak structures on top of a smooth background. Occasionally, the resonances are accompanied by an interference pattern which may lead to irregular line shapes.
The scattering channels will interact also among themselves. In combination with kinematical and structural conditions, e.g., openings of particle emission thresholds or energy-dependent accidental cancellations or enhancements among resonance and background components, the channel-channel interaction amplitudes may produce resonance-like structures in the partial wave cross section which, however, being of purely dynamical origin, will fall out of the systematics expected by symmetries and the related selection rules. Care has to be taken to identify, isolate, and separate model-dependent effects which are achieved by comparing independently derived results of different approaches. Careful studies of the evolution of the amplitudes in the complex plane by Argand diagrams (named after Jean-Robert Argand (1768-1822)) [83] will reveal the nature of the structure. Dalitz plots, introduced in 1953 by Richard Henry Dalitz (1925-2006) while studying kaon decay [84], are indispensable tools for investigations of three-body reaction channels by revealing correlations between decay products.
4 Channel Coupling in Meson-Meson Scattering
The above formalism is a general scheme, applicable to any interacting quantum system at any scale. The same kind of basic mechanisms are acting in baryon-baryon, meson-baryon, and meson-meson systems. Prominent cases are the formation of pion-nucleon resonances like the first excited state of the nucleon given by the well-known Delta-resonance $P_{33}(1232)$ of width $\Gamma_{\Delta}=120~{}\mathrm{MeV}$, corresponding to a lifetime of about $t_{1/2}\sim 10^{-23}$ sec. The widths and finite lifetimes of the members of the $J^{P}=1^{-}$ vector meson octet point in the same direction. The isovector ($I=1$) $\rho(770)$ meson, for example, is embedded into the $\pi\pi$ $P$-wave continuum, resulting in a decay width of $\Gamma_{\rho}\sim 150~{}\mathrm{MeV}$ with a two-pion branching ratio of 99.9%. Isospin selection rules are affecting the rho-meson by excluding the $\rho^{0}\to\pi^{0}\pi^{0}$ decay channel.
Occasionally, the $\rho$ meson is considered as the gauge boson of a broken hidden local chiral symmetry which, however, is distinct from the broken global chiral symmetry of which the pion is the respective Goldstone (named after Jeffrey Goldstone (born 1933)) boson (named after Satyendra Nath Bose) [85]. Moreover, the non-strange vector mesons $\rho,\omega,\phi$ play a central role in Sakurai’s Vector Meson Dominance (VMD) model (named after Jun John Sakurai (1933-1982)) [86] where they account for the (virtually admixed) hadronic part of the photon by which the photon couples to hadronic matter.
The $\rho$- and $\omega$-mesons were detected first in 1961 at LBL. Remarkably, the $\omega(782)$ meson, being the isoscalar partner of the rho-meson, is a rather long-lived object with a small width of about $\Gamma_{\omega}\sim 8~{}\mathrm{MeV}$. The increase in life-time/decrease in decay width by about a factor of 20 is an effect of spin-parity conservation laws, inhibiting decays of an isoscalar vector meson into two pions. An even more extreme case is the $\phi(1020)$ meson which is the singlet partner of the $\omega$-meson. Although located in a mass region with many open channels, a width as small as $\Gamma_{\phi}=4.249\pm 0.013~{}\mathrm{MeV}$ [38] is observed. The long life-time of the $\phi$-meson is caused by the dominant $s\bar{s}$ structure which requires the creation of $u\bar{u}$ or $d\bar{d}$ pairs out of the vacuum before a decay into lighter mesons, e.g., the prevailing $\phi\to K\bar{K}$ process, can occur, as expressed in the OZI rule which is also responsible for extraordinarily small widths of $c\bar{c}$ states like $J/\psi$.
Since presently dedicated meson production facilities are not available, $N+\pi\to N^{\prime}+2\pi$ and $N+\gamma\to N^{\prime}+2\pi$ reactions are used to investigate the otherwise inaccessible $\pi\Delta(1232)$ and $\rho(770)N$ final states, also aiming at separating these final states kinematically (see, for instance, Ref. [87]). Such indirect methods are also applied for extracting hadron-hadron cross sections from hadron production data in high-energy nucleon-nucleus and nucleus-nucleus collisions at the operating hadron facilities.
7 Coupled Channels Models for Baryon Spectroscopy
1 Overview
Modeling hadron production and especially the search for resonances on the quantitative level is a demanding task. Several groups have developed coupled channel theories, derived sophisticated models apt for practical work, and cast the theoretical results into numerical codes, ready for large-scale analyses of the data measured at the experimental sites and, not least, serving also for the preparation of experimental campaigns.
Figure 10: Diagram illustrating the Bethe-Salpeter equation of the T-matrix for scattering of a hadron or photon (dashed line) on a nucleons or any other baryon (full line). The bare interactions, i.e. the Born terms, are denoted by $V$, the scattering amplitude $T$ includes the full scattering series, summed to all orders. The evolution of the intermediate system is described by the Bethe-Salpeter propagator $\mathcal{G}_{BS}$ (see text).
The existing approaches are utilizing a covariant description of the scattering process. The major step is to construct out of an effective Lagrangian the interaction matrix elements [^15] Joseph-Louis Lagrange (1736-1813) worked out the formalism of analytical classical mechanics, and the variational methods which in quantized form became the essential tools of QFT. which serve as input to a system of coupled Bethe-Salpeter equations (named after Hans Albrecht Eduard Bethe (1906-2005) and Edwin Salpeter (1924-2008)) [^16] Hans Bethe received the Nobel Prize in Physics in 1967 for his work on the theory of stellar nucleosynthesis, which is illustrated in Fig. 10. Couplings between channels, e.g., the aforementioned case of $|KN\rangle\leftrightarrow|\pi\Sigma\rangle$ coupling and interference in $\Lambda(1405)$ spectroscopy, have to be taken into account. In Fig. 11, the elemental ingredients of coupled channels approaches for hadron production on the nucleon are displayed in diagrammatical form.
In short-hand notation, the BSE for the transition matrix (T-matrix) reads:
$T(p^{\prime},p;w)=V(p^{\prime},p;w)+\int\frac{\mathrm{d}^{4}q}{(2\pi)^{4}}V(p^{\prime},q;w)\mathcal{G}_{BS}(q;\sqrt{s})T(q,p;w)\>,$ (11)
where $V$ is the matrix containing the full set of elemental channel interactions. $w=\sqrt{s}$ is the available center-of-mass energy. $p$ and $p^{\prime}$ are the incoming (outgoing) hadron (on-shell) four-momenta. Intermediate propagation is described by the BS propagator $\mathcal{G}_{BS}$. The integration over the energy variable $q^{0}$ is performed in closed form by contour integration, leading to the reduced propagator $G_{BS}(\mathbf{q};\sqrt{s})$. That step in fact requires special attention because it involves also a projection to positive energy states, hence effectively eliminating the vacuum contributions. The price to pay is that interactions have to account implicitly for the eliminated degrees of freedom by effective, phenomenological coupling constants. A reduction scheme widely used, e.g., in NN scattering is the Blankenbecler-Sugar (BbS) approach (Richard Blankenbecler (born 1933) and Robert Sugar [88]). The BbS projection reduces the full BSE to a three-dimensional problem by preserving covariance and two- and three-body unitarity.
The solution of a large system of coupled integral equations defined in complex algebra is a formidable numerical task even for modern computing facilities. In order to optimize the numerical workload, the full CC problem is solved piecewise by constructing first the K-matrix and then retrieving the full T-matrix in a second step.
For that goal, the reduced propagator $G_{BS}$ is decomposed into its real and imaginary parts. Accordingly, the (reduced) BSE separates can be rewritten in terms of two (nested) equations. The scattering series generated by $Re(G_{BS})$ can be summed separately, unaffected by $Im(G_{BS})$ leading to the K-matrix equation
$K(\mathbf{p},\mathbf{p}^{\prime};w)=V(\mathbf{p},\mathbf{p}^{\prime})+\int\frac{d^{3}q}{(2\pi)^{3}}V(\mathbf{p},\mathbf{q})\mathrm{Re}\left[G_{BS}(q;w)\right]K(\mathbf{q},\mathbf{p}^{\prime};w)\>,$ (12)
which accounts for all off-shell contributions. The momentum integrals must be regularized, which is accomplished by attaching to the bare interaction $V$ vertex form factors:
$F(q^{2})=\frac{\Lambda_{q}^{4}}{\Lambda_{q}^{4}+(q^{2}-m^{2})^{2}}\>,$ (13)
where $q^{2}$ denotes the four-momentum squared of the involved particle of mass $m$ which mediates the interactions.
The second step consists of reconstructing the complete scattering amplitude by solving the BSE, now given in the form
$T(\mathbf{p},\mathbf{p}^{\prime};w)=K(\mathbf{p},\mathbf{p}^{\prime};w)+i\int\frac{d^{3}q}{(2\pi)^{3}}T(\mathbf{p},\mathbf{q};w)\mathrm{Im}\left[G_{BS}(q;w)\right]K(\mathbf{a},\mathbf{p}^{\prime};w)\>.$ (14)
This is still a system of coupled integral equations but of a much simpler structure than before by the fact that the imaginary part of $G_{BS}$ is given by Dirac $\delta$-distributions in energies and/or powers of momenta, projecting the intermediate channels to their respective on-shell kinematics. The reduced BSE is finally solved by expanding the matrix elements into partial wave components, by which the remaining integration over the angles of the momenta in the chosen reference frame is performed analytically.
Denoting the partial wave components by their total spin $J$, orbital angular momentum $L$ and parity $P$, a further reduced system of coupled integral equations is obtained:
[…]
Final, initial, and intermediate $meson-baryon$ channels are denoted by $f,i,j$, respectively. Spectral distributions of the intermediate configurations are taken into account by integration over the respective mass distribution $A_{j}(\mu_{j}$, thus allowing a detailed description of unstable states like $f_{0}\sigma(500),\rho(770)\ldots$ attached to a stable baryon or the propagation and interaction of decaying baryons like $\Delta_{33}(1232)$
For stable particles of mass $m_{j}$, the mass distributions reduce to $A_{j}=\delta(\mu_{j}-m_{j})$. In the other cases, the integrals are evaluated by numerical integration formulas, thus replacing the integrals by a sum over a discrete set of mesh points. Then, the integral equations reduce to a system of coupled linear equations that are solved numerically. The essence of such a treatment is that unstable states are represented by a discrete distribution of states with fractional spectroscopic strengths as defined by the value of the spectral distribution at the mesh points and the integration weights, see, e.g.[89]. Unitarity is fulfilled as long as $V$ is Hermitian.
2 Perturbative Treatments: Photon-Hadron Channels and $K$-matrix Born Approximation
The much weaker electromagnetic coupling constant, $e^{2}\sim\alpha_{f}\simeq 1/137\ll g^{2}_{\pi N}\sim 14$, the photo-production reaction channels may be treated perturbatively in leading order of the $\gamma N$, $\gamma N^{\ast}$, and photon-hadron vertices in general. Thus, the T-matrix elements involving a photon are treated in the lowest order Born approximation, $T_{\gamma h}\to V_{\gamma h}$, where $h$ denotes a hadron.
A significant reduction of the numerical effort is achieved by the so-called $K$-matrix Born approximation by using $K\equiv V$ [90]. This approximation corresponds to neglecting the real part of $G_{BS}$. Hence, it is in fact a pole approximation since the intermediate propagation is frozen to the on-shell contributions produced by the remaining imaginary part of $G_{BS}$.
The systematic term-by-term summation of scattering series and similar perturbative series goes back to Max Born (1882-1970). Max Born shared the 1954 Nobel Prize in Physics with Walther Bothe (1891-1957) for his fundamental research in quantum mechanics, especially in the statistical interpretation of the wave function. Moreover, Max Born’s book of 1933 on Optik [91] played a central role in theory and experiments on gamma spectroscopy. In 1959, the book was updated, extended by contributions from other authors, and republished in the English language, experiencing several reprints [92].
3 Interaction Potential and Scattering Matrix
The hadron-hadron interaction matrix $V$ is built by a sum of $s$-, $u$-, and $t$-channel Feynman diagrams (once invented by Richard Feynman for visualizing QFT/QED amplitudes) of the type shown in Fig. 11. They are derived from the underlying effective model Lagrangians, e.g., see [93] for the one used in GiM. In order to reduce the number of model parameters, the non-resonant background terms should be derived consistently by the $u$- and $t$-channel diagrams resulting from the model Lagrangian.
Figure 11: Structure of the tree-level interaction potential $V$. $s-$, $u-$, and $t-$ channel interactions defining the non-resonant background contributions are shown in the first line, including contact terms which are chosen such that gauge invariance is assured. The so-called z-diagrams, displayed in the second line are generic for the double-pion channels. $s-$ channel resonance interactions are depicted in the last line. Time is running from left to right.
In a symbolical short-hand notation, the bare, Born-term interaction for a reaction $a\to b$ at center-of-mass energy $w=\sqrt{s}$ is written as
$V_{ab}=V^{(sut)}_{ab}+V^{(Z)}_{ab}+V^{(R)}_{ab}\>,$ (16)
where the contributions represent the first, the second, and the third line of Fig.11, respectively, where
$V^{(R)}_{ab}=\sum_{Q}\frac{V_{aQ}V_{Qb}}{w-m_{Q}}\>.$ (17)
In each partial wave channel $[JLP]$ a hadronic interaction matrix is found
$\mathcal{V}^{[JLP]}=\left(\begin{array}[]{ccccc}V_{\pi\pi}&V_{\pi\rho}&V_{\pi\sigma}&V_{\pi K}&\ldots\\ V_{\rho\pi}&V_{\rho\rho}&V_{\rho\sigma}&V_{\rho K}&\ldots\\ V_{\sigma\pi}&V_{\sigma\rho}&V_{\sigma\sigma}&V_{\sigma K}&\ldots\\ V_{K\pi}&V_{K\rho}&V_{K\sigma}&V_{KK}&\ldots\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ \end{array}\right)^{[JLP]}$ (18)
where $V_{\pi\pi}\equiv V_{\pi B\pi B^{\prime}}$ for baryons $B,B^{\prime}=N,N^{*},\Lambda,\Sigma\ldots$, and correspondingly for the other meson channels..
After deriving the K-matrix, the hadronic partial wave T-matrices are defined by a system of coupled linear equations with the formal solution
$\mathcal{T}^{[JLP]}=\left(1-i\rho_{C}\mathcal{K}^{[JLP]}\right)^{-1}\mathcal{K}^{[JLP]}\approx\left(1-i\rho_{C}\mathcal{V}^{[JLP]}\right)^{-1}\mathcal{V}^{[JLP]}\>,$ (19)
where also the K-matrix Born approximation is indicated. The phase space factor $\rho_{C}$ is produced by evaluating the integral over the delta-distributions of $Im[G_{BS}]$.
Adding the perturbatively determined photo-production T-matrix elements, the complete T-matrix is finally obtained
$\mathcal{T}^{[JLP]}=\left(\begin{array}[]{cccccc}T_{\gamma\gamma}&T_{\gamma\pi}&T_{\gamma\rho}&T_{\gamma\sigma}&T_{\gamma K}&\ldots\\ T_{\pi\gamma}&T_{\pi\pi}&T_{\pi\rho}&T_{\pi\sigma}&T_{\pi K}&\ldots\\ V_{\rho\gamma}&T_{\rho\pi}&T_{\rho\rho}&T_{\rho\sigma}&T_{\rho K}&\ldots\\ V_{\sigma\gamma}&T_{\sigma\pi}&T_{\sigma\rho}&T_{\sigma\sigma}&T_{\sigma K}&\ldots\\ V_{K\gamma}&T_{K\pi}&T_{K\rho}&T_{K\sigma}&T_{KK}&\ldots\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ \end{array}\right)^{[JLP]}.$ (20)
4 Couple Channels Projects for Partial-Wave Analyses
Historically, Adrien-Marie Legendre (1752-1833) was the first to find the multipole expansion of the gravitational two-body potential [94] for which he had to invent his famous system of orthogonal polynomials, since then indispensable for any partial wave expansion. Rayleigh’s book on the Theory of Sound [95] established PWA as a tool in classical wave mechanics. In quantum scattering theory PWA is the mathematical tool to calculate reaction amplitudes for given interactions. On the practitioner’s side PWA is the key technique for determining reaction amplitudes by fits to scattering data and concluding on the underlying interaction. That task is in fact a non-trivial mathematical problem, corresponding to the solution of an in general ill-posed inversion problem - as was pointed out some time ago on quantum-mechanical grounds, e.g., by Andrei Nikolaevich Tikhonov [96] (1906-1993).
Over the years, in hadron physics several CC approaches have been developed, out of which a few long-term projects have emerged. The investigations of Höhler’s [97] and Cutkosky’s group [98] have the merits of establishing computational methods in hadron spectroscopy as tools for systematically enlarging and interpreting the database. Coupled-channel (CC) approaches have been proven to be an efficient tool to extract hadron properties from experiments. They are firmly established as indispensable workhorses for research on hadron spectroscopy, playing the dominant role, especially in baryon spectroscopy.
In that tradition, the SAID project is probably the one with the largest long-term impact on the community. SAID, initiated by Richard A. Arndt, L. David Roper, and their group at Virginia Polytechnic Institute and State University and since then fostered continuously at The George Washington University since 1999 [99], provides compilations of experimental data, constant updates of CC methods and results. The easy online accessibility is an important service to the community and strongly supports the research on hadron spectroscopy. Unlike the other CC approaches referred to below, SAID does not assume ad-hoc resonance contributions but derives partial wave amplitudes directly from data, followed by an analysis on pole positions in the complex energy plane and BW parameters. The latter approach has natural limits because of: (i) revealing wide resonances up to a width $\Gamma<500~{}\mathrm{MeV}$, (ii) missing transitions of small Branching Ratios $BR<4\%$, and (iii) tending (by construction) to miss narrow resonances if $\Gamma<20~{}\mathrm{MeV}$. However, in SAID narrow resonances are accounted for by a modified PWA method [100, 101]. The computational SAID approach accounts for energy dependencies of self-energies on the basis of the Chew-Mandelstam method.
The MAID project, connected to the MAMI facility at Johannes Gutenberg University of Mainz, has been following similar routes. The Bonn-Gatchina (BnGa) PWA project of Bonn University and Saint Petersburg Nuclear Physics Institute (PNPI) at Gatchina [^17] Gatchina is a town in $40~{}\mathrm{km}$ on the south of Saint Petersburg and by about $700~{}\mathrm{km}$ far from Moscow. collaboration and centered at the ELSA laboratory at Bonn University is another influential CC activity [102]. BnGa utilizes the so-called N/D approach (see [103]) and otherwise incorporates loop corrections into the K-matrix approach in a spirit similar to SAID. The BnGa, MAID, and the SAID solutions are accessible online.
The Jülich-Bonn-Washington Collaboration (Jü-Bo), including Jülich Center für Hadron Physics, Bonn University, and The George Washington University, has been formulating a coupled channels model [104, 105]. That approach, in fact, is based largely on the work of Haberzettl et al. [106].
Over a long period, the Giessen group at Justus Liebig-University of Giessen has been developing its own covariant CC model, the Giessen Model (GiM). Initiated in the early 1990s by Ulrich Mosel, the project has grown over a few decades to a full-fledged and versatile tool for baryon spectroscopy by meson photoproduction on the nucleon and subsequent spectroscopy of the decay channels. The GiM utilizes a K-matrix approach strictly oriented on Lagrangian methods with special attention on the consistency of interactions [107]. The theoretical background of GiM is closely oriented to the discussions in the previous paragraphs. Applications range from single pion, eta, and kaon production and investigations of vector meson production to the population of two-pion channels and Compton scattering [108] [^18] Named after Arthur Holly Compton (1892-1962) who won the 1927 Nobel Prize in Physics for his discovery of the Compton effect, which demonstrated the particle nature of electromagnetic radiation.. A recent overview of the GiM is found in [109].
5 Quantum Interference in Hadron Spectra
At the time of writing this article, the science and technology world is celebrating 100 years of quantum mechanics. Among others, coherence of wave functions and interference of matrix elements are defining properties of quantum systems. While in atomic, molecular, and solid state systems quantum interference is used actively to manipulate the systems, see, e.g.[110], the vastly different conditions encountered in nuclear and sub-nuclear systems usually inhibit direct actions at a comparable level. Important signatures for interference phenomena are irregularities in line shapes, observed as significant deviations from Lorentz (named after Hendrik Antoon Lorentz (1853-1928) [^19] Lorentz shared the 1902 Nobel Prize in Physics with Pieter Zeeman (1865-1943) for their discovery and theoretical explanation of the Zeeman effect (named after Pieter Zeeman).) (or Gaussian (named after Carl Friedrich Gauss (1777-1855)) shapes as expected for isolated, non-interacting resonances. In nuclear spectra, exceptional line shapes have been seen in a few cases [111, 112, 113] and interpreted in the Fano formalism [^20] Ugo Fano (1912-2001), originally a member of Fermi’s team, was a master at understanding how radiation interacts with matter. His work set the agenda for much of modern atomic physics, and had a broad sweep across field as stated in an obituary published in [114]. [115], being used also in atomic and laser physics in [110]. In [116, 117] that formalism was applied to interpret irregular line shapes of the $D\bar{D}$ decay modes of the aforementioned $\psi(3770)$ resonance.
Though the phenomenon of quantum-mechanical interference has been known for many years, there are still open questions, not only with respect to hadron physics but also in electronic systems, see, e.g.[118]. In the review [119], Azimov discussed how the interference of resonances may and does work. A rich source is data on rare decay modes of well-known resonances which demonstrate a wide variety of possible different manifestations of interference. Some special kinds of resonance interference, not yet sufficiently studied and understood, are also briefly considered. The interference may give useful experimental procedures to search for new resonances with arbitrary quantum numbers, even with exotic ones, and to investigate their properties.
In Figure 12 resonance interference is illustrated for $\rho^{0}$, $\omega$-, and $\phi$-meson production in the reaction $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}$. Total cross sections collected by the SND Collaboration are shown. Actually, for the $3\pi$ decay case, $\Gamma(\rho^{0}\to 3\pi)=0.015~{}\mathrm{MeV}$ (Isospin symmetry violation), $\Gamma(\omega\to 3\pi)=7.58~{}\mathrm{MeV}$, and $\Gamma(\phi\to 3\pi)=0.65~{}\mathrm{MeV}$ (Zweig rule violation) [38]. It has a clear BW-like peak in the ($\rho^{0}$, $\omega$) region and a bump–dip structure in the $\phi$ region. Background near $\phi$ changes slowly vs. nearly standard interference curve. Instead of $\phi$–peak, there are both bump and dip, each has a form different from BW; max/min shifts from $\phi$-mass. There is a similar shift for the $\rho$-mass. Then $\rho$- contribution here deforms the $\omega$-tail.
Figure 12: The $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}$ cross section measured by the SND Collaboration and collected in [120]. The curve is the fit with the $\rho^{0}$, $\omega$, $\phi$, $\omega^{\prime}$, and $\omega^{\prime\prime}$ resonances. Red arrows shown $\rho^{0}$, $\omega$, and $\phi$ thresholds.
The interpretation of the recent BESIII cross sections for the reaction $e^{+}e^{-}\to J/\psi\pi^{+}\pi^{-}$ [121] as an observation of the $X(3872)$ resonance [122] via a destructive interference between small resonance and large background contributions (that is a natural way for the inelastic reaction, see, for instance, [123]) (Fig. 13). The large non-resonance contribution magnifies a small resonance effect, as described in different contexts by the Fano mechanism. The observed dip is small, but visible, and more statistics would be highly desirable.
Figure 13: The line shapes for the three best fits [122] to the BESIII data [121] for the reaction $e^{+}e^{-}\to J/\psi\pi^{+}\pi^{-}$ after convolution with the energy spread function. As an example, the blue dashed line shows the line shape for fit 1 without the effect of the energy spread. The $1\sigma$ error bands correspond to the uncertainty propagated from the data.
The data displayed in Figs. 12 and 13 demonstrate the direct interference between two resonances decaying into the same final states [^21] A while ago Richard Feynman said: When looking at Maxwell equations, it is hard to imagine how beautiful the rainbow is. Then Yakov Azimov (1938-2016) extended it: Everybody knows that the interference does exist. But it is not always easy to imagine how it will work in a particular case..
8 Meson Photoproduction
The determination of the resonance properties for all accessible baryon states is a central objective in nuclear physics. The extracted resonance parameters provide a crucial body of data for understanding the dynamics of nucleon excitations and the spectral distributions, essential for testing o´phenomenological models of the nucleon, calibrating lattice QCD calculations as a way of connecting non-perturbative hadron physics to QCD. The spectrum of $N^{\ast}$ and $\Delta^{\ast}$ non-strange baryon resonances of masses up to about 2 GeV is probably the best-studied sector of hadron physics. Meson-nucleon scattering and photoproduction of mesons on nucleon and nuclear targets have led to a wealth of data, lately also supplemented by combined systematic studies of kaon and associated hyperon production experiments. Properties of the known resonances continue to become better determined as experiments involving polarized beams, targets, and recoil measurements are expanded and refined. With increasing energy, multi-meson channels will open. They provide the opportunities to identify states coupled weakly to two-body channels like $\pi+N$. Since production reactions involving the octet - neglecting the $\eta^{\prime}$ singlet - of pseudo-scalar mesons and the related baryon resonances are studied the best, they will be in the center of the following discussions. To a lesser degree, also production reactions leading to the nonet of vector mesons have been investigated and will be addressed selectively. Thus, in the following sections, reactions involving the baryon octet and the decuplet, Fig. 6 and two of the three meson nonets of Fig. 7 will be discussed.
The scalar mesons are suffering from short lifetimes and ill-defined spectral distributions, see, e.g., the status of the $f_{0}(500)/\sigma$ meson [38]. Moreover, quark model studies predict that the scalar mesons of masses below $1~{}\mathrm{GeV}$ are most likely tetra-quark states of $qq-\bar{q}\bar{q}$ structure [124]. Therefore, light scalar meson production on the nucleon will not be considered. Accordingly, reactions with direct multi-meson production will be neglected, too.
1 Single Pseudoscalar Meson Photoproduction
Pion Photoproduction on Proton
Single pseudoscalar meson photoproduction involves the interaction of a photon with a free proton, a bound neutron, or a whole nucleus. For studies of the baryon spectrum, one is normally interested in the first two of these. So a spin-1 particle (the photon, two helicity states) and a spin-$\frac{1}{2}$ particle (the nucleon) react to give a spin-0 particle (the pseudoscalar meson) and a spin-$\frac{1}{2}$ particle (the recoiling baryon). This gives eight spin combinations, of which four are possible within the parity-conserving strong interaction that has taken place. EM interaction does not conserve isospin, so multipole amplitudes contain isoscalar and isovector contributions of EM current.
The four combinations are represented as amplitudes (Eqs. (21) and (22)), the exact form of which is a matter of choice. A theory of pion photoproduction was constructed in the 1950s. Kroll and Ruderman [125] were the first to derive model-independent predictions in the threshold region, a so-called low energy theorem (LET), by applying gauge and Lorentz invariance to the reaction $\gamma N\to\pi N$. The general formalism for this process was developed by Chew, Goldberger, Low, and Nambu (Geoffrey Foucar Chew (1924-2018), Marvin Leonard “Murph” Goldberger (1922-2014), Francis Eugene Low (1921-2007), and Yoichiro Nambu (1921-2015) [^22] Nambu was awarded half of the Nobel Prize in Physics in 2008 for the discovery in 1960 of the mechanism of spontaneous broken symmetry in subatomic physics, related at first to the strong interaction’s chiral symmetry and later to the electroweak interaction and Higgs mechanism.). The results are known as CGLN amplitudes [126] and a few years later helicity amplitudes [127] were derived. About a decade later, Berends, Donnachie, and Weaver analyzed the existing data in terms of a multipole decomposition and extracted the various multipole amplitudes contributing in a region up to an excitation energy of $500~{}\mathrm{MeV}$ [128]. These amplitudes are vital inputs to low-energy descriptions of hadron physics based on the chiral perturbation theory (ChPT) [129]. Within any of these bases, there are 16 possible bilinear combinations that are referred to as the “observables.”
For the proton
$A(\gamma p\to\pi^{0}p)=A^{(0)}+\frac{1}{3}A^{(1/2)}+\frac{2}{3}A^{(3/2)}~{}~{}~{}\mathrm{and}~{}~{}~{}A(\gamma p\to\pi^{+}n)=\sqrt{2}\biggl{(}A^{(0)}+\frac{1}{3}A^{(1/2)}-\frac{1}{3}A^{(3/2)}\biggl{)}\>.$ (21)
For the neutron
$A(\gamma n\to\pi^{0}n)=\sqrt{2}\biggl{(}-A^{(0)}+\frac{1}{3}A^{(1/2)}+\frac{2}{3}A^{(3/2)}\biggl{)}~{}~{}~{}\mathrm{and}~{}~{}~{}A(\gamma n\to\pi^{0}n)=-A^{(0)}+\frac{1}{3}A^{(1/2)}+\frac{2}{3}A^{(3/2)}\>.$ (22)
Proton data alone does not allow separation of isoscalar and isovector components [130].
Accurate evaluation of EM couplings $N^{\ast}\to\gamma N$ and $\Delta\to\gamma N$ from meson photoproduction data remains a paramount task in hadron physics. Only with good data on both proton and neutron targets can one hope to disentangle isoscalar and isovector EM couplings of various $N^{\ast}$ and $\Delta^{\ast}$ resonances [131, 132], as well as isospin properties of non-resonant background amplitudes.
The lack of $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$ data [^23] Apart from lower-energy ($<700~{}\mathrm{MeV}$), there are data for the inverse $\pi^{-}$ photoproduction reaction, $\pi^{-}p\to\gamma n$ [99]. This process is free from complications associated with a deuteron target. However, there is a major disadvantage of using $\pi^{-}p\to\gamma n$: there is a large background from $\pi^{-}p\to\pi^{0}n\to\gamma\gamma n$, whose cross section is 5 to 500 times larger than $\pi^{-}p\to\gamma n$. does not allow us to be as confident about the determination of neutron couplings relative to those of protons.
Since the neutron targets do not exist, it remains to use nuclear ones. In this case, when extracting information on the elementary reaction on the bound neutron from nuclear data, one should take into account the nuclear-medium effects, i.e., the final-state interaction (FSI) and Fermi-motion effects [133, 134] [^24] It is impossible to measure FSI experimentally [135]. Obviously, in the case of polarized measurements the FSI corrections are small or consistent with experimental uncertainties.
Measurements of pion photoproduction on both proton and quasi-free neutron targets have a very long history, starting about 70 years ago with the discovery of the pion by the University of Bristol group [4]. Two years later, at the 1949 Spring Meeting of the US National Academy of Sciences, a preliminary account was given of some observations of mesons produced by the $335-\mathrm{MeV}$ photon beam from the Berkeley synchrotron [136] [^25] Edwin Mattison McMillan (1907-1991) shared the 1951 Nobel Prize in Chemistry with Glenn Theodore Seaborg (1912-1999), credited with being the first to produce a transuranium element, neptunium. In addition, McMillan co- invented the synchrotron with Vladimir Iosifovich Veksler (1907-1966). Starting with the use of bremsstrahlung facilities, pioneering results for $\gamma p\to\pi^{+}n$ (Fig. 14[136] and for $\gamma p\to\pi^{0}p$ [137] were obtained. One can possibly understand that the pion family is a triplet. Finding a neutral pion is much more difficult since it does not leave marks in photoemulsion or Wilson chambers (named after Charles Wilson).
In addition to the free proton reactions as $\gamma p\to n\pi^{+}$, the experiment at Berkeley has utilized the loosely bound neutron $\gamma n\to p\pi^{-}$ at the photon energy $318\pm 10~{}\mathrm{MeV}$ [138]. So, the negative pion photoproduction experiment was done in the next 3 years after the positive one.
Figure 14: Distribution of positive pion energies from photon energy of $335~{}\mathrm{MeV}$ [136]. The apparent lower limit on the energy is caused by the fact that the energies are computed as if the mesons originated in the center of the carbon block. The dashed line is simply a guess as to the trend of the distribution at low energies, which was used in the integration leading to the total cross section.
Despite all the shortcomings of the first measurements (such as large normalization uncertainties, wide energy and angular binning, limited angular coverage, and so on), these data were crucial for the discovery of the first excited Delta and nucleon states using PWA for $\pi N$ elastic scattering data. The $\Delta(1232)3/2^{+}$ was determined by Enrico Fermi’s group [139]. While the second one $N(1440)1/2^{+}$, called “Roper,” (named after L. David Roper (born 1935) who discovered this baryonic state) came in several years after the $\Delta$-isobar [140] [^26] As an anecdote, it is worth mentioning that Lev Davidovich Landau, when becoming aware of Fermi’s discovery of the Delta resonance, did not believe in the existence of such a state, arguing A width of $120~{}\mathrm{MeV}$ what is that? The pion will make a quarter of the circle around the nucleon and that is supposed to be a pion-nucleon bound state? The concept of a meson-nucleon resonance was waiting to be established $\ldots$..
Figure 15: Database for $\gamma p\to\pi^{0}p$ (left) and $\gamma p\to\pi^{+}n$ (right). Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
Figure 16: Database for $\gamma n\to\pi^{-}p$ (left) and $\gamma n\to\pi^{0}n$ (right). Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
Pion Photoproduction on the Neutron
The “neutron” database is significantly smaller than the “proton” one. The majority of single pseudoscalar meson photoproduction of the “neutron” target came from EM facilities at BNL and JLab, USA; MAMI, Germany; GRAAL, France (Fig. 16[141].
Studies of the $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$ reactions can be carried out in quasi-free kinematics with deuteron targets. The reactions $\gamma d\to\pi^{-}p(p)$ and $\gamma d\to\pi^{0}n(p)$ in these kinematics have a fast, knocked-out nucleon and a slow proton spectator, and the slow proton is assumed not to be involved in the pion production process. In this quasi-free region, the reaction mechanism corresponds to the “dominant” impulse approximation (IA) diagram in Fig. 17(a) with the slow proton emerging from the deuteron vertex. Here, the differential cross section on the deuteron can be related to that on the neutron target in a well-understood way, see Ref. [142] and references therein. Fig. 17 illustrates this dominant IA diagram, as well as the leading terms of FSI corrections.
An energy and angle dependent FSI correction factor, R(E, $\theta$), can be defined as the ratio between the sum of three dominant diagrams in Fig. 17 and IA (the first of the diagrams). This can then be applied to the experimental $\gamma d$ data to get a two-body cross section for $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$.
The GWU-ITEP FSI calculations (see Ref. [142] and references therein) are available over a broad energy range (threshold to $E=2.7~{}\mathrm{GeV}$), and for the full CM angular range ($\theta=0^{\circ}$ to $180^{\circ}$). Overall, the FSI correction factor $R<1.0$, while its value varies from 0.70 to 0.90 depending on the kinematics. The behavior of R is very smooth vs. pion production angle. There is a sizable FSI effect from the $S$-wave part of $pp$-FSI at small angles.
Figure 17: Feynman diagrams for the leading components of the $\gamma D\to\pi NN$ amplitude. a: Impulse approximation (IA), b: $NN$-FSI, and c: $\pi N$-FSI. Filled red circles show FSI vertices. Wavy, dashed, solid, and double lines correspond to the photons, pions, nucleons, and deuterons, respectively.
R(E, $\theta$) is used as the FSI correction factor for the CLAS quasi-free $\gamma d\to\pi pN$ cross section averaged over the laboratory photon energy bin width. Note that the FSI correction grows rapidly to the forward direction ($\theta<30^{\circ}$). There are currently few measurements in this regime, so the uncertainty due to FSI for this reaction at forward angles does not cause too much concern. The contribution of uncertainty in FSI calculations to the overall systematic normalization uncertainty is estimated to be about 2%–3% (the sensitivity to the deuteron wave function is 1% and to the number of steps in the integration of the five-fold integrals is 2%). For the CLAS measurements, no sensitivity was found to the value of proton momentum used to determine whether or not it is a spectator.
The $\gamma n\to\pi^{0}n$ measurement is much more complicated than the case of $\gamma n\to\pi^{-}p$ because the $\pi^{0}$ can come from both neutron and proton initial states. The GW-ITEP studies have shown that photoproduction cross sections from protons and neutrons are generally not equal [135]. For $\pi^{0}$ photoproduction on proton and neutron targets, one has
$A(\gamma p\to\pi^{0}p)=A_{v}+A_{s}~{}~{}~{}\mathrm{and}~{}~{}~{}A(\gamma n\to\pi^{0}n)=A_{v}-A_{s}\>,$ (23)
where $A_{v}$ and $A_{s}$ are the isovector and isoscalar amplitudes, respectively. Therefore, if $A_{s}\neq 0$ then the $\gamma p$ and $\gamma n$ amplitudes are not equal.
Eta and Eta’ Photoproduction on the Proton
The experimental activity on the pseudoscalar $\eta^{(}548)$ and $\eta^{\prime}(958)$ mesons has a short history, not to the least because they are charge-neutral particles and, as such, are notoriously hard to detect. The singlet $eta^{\prime}$ meson, in addition, is much heavier than the octet part of the pseudoscalar nonet. Compared to the pion case, the data base is much smaller, as is evident from (Fig. 18).
Figure 18: Database for $\gamma p\to\eta p$ (left) and $\gamma p\to\eta^{\prime}p$ (right). Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
Figure 19: Database for the Kaon photoproduction data on both proton and neutron targets. Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
Kaon Photoproduction
As mentioned before, kaon physics was the primer for realizing that CP symmetry is violated, with far-reaching implications for elementary particle physics. The existence of two separate kaon decay branches was the first definite signal of competing and interfering processes. Moreover, they enforced the concept of superpositions for elementary particles by the necessity to introduce the $K_{L}$ and $K_{S}$ states of the $K^{0}/\bar{K}^{0}$ system. The kaon database is shown in Fig. 19.
Vector Meson Photoproduction on the Proton
Figure 20: Database for $\gamma p\to\omega p$. Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
There are no vector meson beams, so experiments using modern electromagnetic (EM) facilities attempt to access vector meson nucleon to vector meson nucleon interaction via EM production reactions $ep\to e^{\prime}\text{(vector meson)}~{}p$. Statistics for vector meson photoproduction are very limited. There is a map just for the $\omega$ case (Fig. 20). Data came from JLab, MAMI, ELSA, and GRAAL.
Some vector mesons can, compared to other mesons, be measured with very high precision. This comes from the fact that vector mesons have the same quantum numbers as the photon: $I^{G}(J^{PC})=0^{-}(1^{-~{}-})$. It allows us to apply a Sakurai’s Vector Meson Dominance (VMD) model, assuming that a real photon can fluctuate into a virtual vector meson, which subsequently scatters off a target nucleon (Fig. 21[86].
Figure 21: Schematic diagrams of vector-meson photoproduction (left) and the VMD model (right) in the energy region at threshold experiments. $V$ means vector meson.
High-statistics total cross sections for the vector meson photoproduction at the threshold: $\gamma p\to\omega p$ (from A2 at MAMI [143], ELPH [144], and CBELSA/TAPS) [145], $\gamma p\to\phi p$ (from CLAS [146] and LEPS [147, 148]), and $\gamma p\to J/\psi p$ (from GlueX) [149] allow one to extract the absolute value of the vector meson nucleon scattering length (SL) using the VMD model. The extended analysis of $\Upsilon$-meson photoproduction using quasi-data from the QCD approach is in perfect agreement with the light-meson findings using experimental data [150]
Let us focus on four vector mesons ($\omega$, $\phi$, $J/\psi(1S)$, and $\Upsilon(1S)$) from $q\bar{q}$ nonet, the widths of which are narrow enough to study meson photoproduction at threshold and where data and quasi-data are available (Table 2). To avoid a broad width problem at threshold, one is not considering the $\rho$-meson case to determine vector meson nucleon SL. Furthermore, one will ignore, for example, $\psi^{\prime}(2S)$ due to the difference between the $1S$ and $2S$ states due to “zero” in radial wave functions (WFs). Unfortunately, one cannot go above Quarkonium or $\Upsilon$, whose quark content is $b\bar{b}$. The problem is that actually the Toponium ($T(1S)$), whose quark content is $t\bar{t}$, does not exist. It is due to a large mass of the Theta meson and the $t$-quark decays faster than the quarks form the Theta meson.
Table 2: List of the vector mesons including quark contents and width of them [38]. The four vector mesons, used for the determination of the scattering length as discussed in the text, are marked in blue.
Meson Quark $\Gamma$
Content [MeV]
$\rho^{+}(770)$ $u\bar{d}$ 148
$\rho^{0}(770)$ $\frac{(u\bar{u}-d\bar{d})}{\sqrt{2}}$ 149
$\omega(782)$ $\frac{(u\bar{u}+d\bar{d})}{\sqrt{2}}$ 8.5
$K^{\ast+}(892)$ $u\bar{s}$ 51
$K^{\ast 0}(892)$ $d\bar{s}$ 47
$\phi(1020)$ $s\bar{s}$ 4.3
$D^{\ast+}(2010)$ $c\bar{d}$ 0.083
$D^{\ast 0}(2007)$ $c\bar{u}$ $<$2.1
$J/\psi(1S)(3097)$ $c\bar{c}$ 0.093
$\psi^{\prime}(2S)(3686)$ $c\bar{c}$ 0.284
$\Upsilon(1S)(9460)$ $b\bar{b}$ 0.052
Figure 22: Comparison of the $|\alpha_{Vp}|$ SLs estimated from threshold vector meson photoproduction on the proton target with VMD model contribution vs the inverse mass of the vector mesons. Input data for phenomenological analyses came from A2 at MAMI (magenta filled circle) [143], ELPH (black filled triangle) [144], and CBELSA/TAPS (black open square) [145] Collaborations for the $\omega$-meson; CLAS (blue filled diamond) [146] and LEPS (green filled square) [147, 148] Collaborations for the $\phi$-meson; and GlueX (red filled star) [149] Collaboration for the $J/\psi$-meson; and quasi-data from Center for Nuclear Femtography (brown open star) [150] for the $\Upsilon$-meson. Analyses results for $\omega$-meson is given at Refs. [143, 144, 151]; for $\phi$-meson is given at Refs. [152, 151]; for $J/\psi$-meson is given at Ref. [153]; and for $\Upsilon$-meson is given at Ref. [154]. The black dashed line is hypothetical following $|\alpha_{Vp}|\propto 1/m_{V}$. $V$ means vector meson.
The “young” vector meson hypothesis may explain the fact that the obtained SL value for the nucleon $\phi$-meson compared to the typical hadron size of approximately $\sim 1~{}\mathrm{fm}$ indicates that the proton is more transparent for the $\phi$-meson compared to the $\omega$-meson and is much less transparent than the $J/\psi$-meson.
[…]
In a recent study, the effect of the VMD assumption was studied in the formalism of Dyson-Schwinger equations, which one can consider as an alternative interpretation of the “young age” effect in another (more formal) language [156].
The suggested approach can be employed to evaluate the $J/\psi$-nucleon SLs, replacing the photon by a $J/\psi$-meson (Fig. 21). The results then appear to have the order of several units of $10^{-3}~{}\mathrm{fm}$, $\alpha_{J\psi N}^{(J=1/2)}=(0.2...3.1)\times 10^{-3}~{}\mathrm{fm}$ and $\alpha_{J\psi N}^{(J=3/2)}=(0.2...3.0)\times 10^{-3}~{}\mathrm{fm}$, where $J$ corresponds to the total angular momentum of the $J/\psi$-nucleon system [157].
Future high-quality experiments by EIC and EicC will have the opportunity to evaluate cases for $J/\psi$- and $\Upsilon$-mesons. It allows one to understand the dynamics of $c\bar{c}$ and $b\bar{b}$ production at the threshold. The ability of J-PARC to measure $\pi^{-}p\to\phi n$ and $\pi^{-}p\to J/\psi n$, which are free from the VMD model, is considered.
2 Single Pseudoscalar Meson Electroproduction
An important spectroscopic tool, complementary to photoproduction, is electroproduction of mesons on the nucleon. As indicated in Fig. 23 electron-induced meson production proceeds by the exchange of a virtual photon which allows for the transfer of energy and momentum beyond the on-shell constraints inherent to photoproduction.
Figure 23: Feynman diagram illustrating electroproduction of a $\pi^{+}$ meson on a proton by an incident electron through exchange of a virtual (off-shell) photon $\gamma^{*}$ transferring the off-shell four-momentum $q=(\omega,\mathbf{q})^{T}$.
Thus, electroproduction data are an important tool to study the properties of non-strange baryons simultaneously in energy and three-momentum transfer as independent variables. Therefore, ongoing PWA fits incorporate the available electroproduction data. The map of $Q^{2}$ dependence of the pion electroproduction data ($\gamma^{\ast}p\to\pi^{0}p$, $\gamma^{\ast}p\to\pi^{+}n$, and $\gamma^{\ast}n\to\pi^{-}p$, no data for $\gamma^{\ast}n\to\pi^{0}n$) before 2009 is shown in Fig. 24. One notes that the CLAS Collaboration produced more than 85% of the world’s pion electroproduction data (this database is still growing), much of which was focused on the mapping of the properties of the $\Delta(1232)$ and higher resonances. Useful comparisons, for instance, will require those involved in this effort to make available all amplitudes obtained in any new determination of ratios $R_{EM}$ and $R_{SM}$ for the transition $N\to\Delta(1232)$ which may be compared with LQCD calculations. Their values are far from those expected of the perturbative regime, $R_{EM}=1$, and $R_{SM}$ is $Q^{2}$ independent [158].
Figure 24: $Q^{2}$ distribution of pion electroproduction data which are now available [159]. Numbers in middle each plot corresponded to number of events. Red filled circles show the number of pion photoproduction events.
3 Strangeness Production on the Nucleon
Strangeness production on the nucleon by excitation of resonances which decay into kaon-hyperon channels is an important spectroscopic tool giving access to the SU(3) flavor structure of baryons. Moreover, such exotic channels like the kaon-hyperon final states are expected to play a key role in identifying hitherto undetected excited states of the nucleon, thus addressing the notorious problem of missing resonances. In [160], pion- and photon-induced $K\Lambda$ reactions were studied within the unitary coupled-channel effective Lagrangian approach. Data on the photoproduction of kaons on the nucleon from the SAPHIR, CLAS and the CBELSA experiment were described by the GiM coupled channels K-matrix approach, also taking into account the full set of all other meson-baryon channels. Thus, a major revision of the complete parameter set was performed. A major goal of those investigations was to address the at that time still open question on the major contributions to the associated strangeness production channels. Since $K\Lambda$ photoproduction data [161, 162] gave an indication for missing resonance contributions, a combined analysis of the $\pi+N\to K\Lambda$ and the $\gamma+N\to K\Lambda$ reactions was expected to identify clearly these states. Assuming small couplings to $\pi N$, these hidden states should not exhibit themselves in the pion-induced reactions and, consequently, in the $\pi N\to K\Lambda$ reaction. The aim of our calculations was to explore to what extent the data available at that time can be explained by known reaction mechanisms without introducing new resonances. Our results for total cross sections are displayed in Fig. 25 and further results on differential cross sections, polarization observables, and angular distributions are found in [160]. As discussed in [160], the SAPHIR [161] and the CLAS [162] data sets, in fact, are leading to two slightly different sets of interaction parameters, reflecting and emphasizing the differences among the two measurements. Below, that point is discussed again.
Figure 25: $\pi^{-}p\to K^{0}\Lambda$ total partial wave cross sections, predicted by parameter set C of Ref. [160], obtained from a fit to the CLAS data [162]. The experimental cross section data are taken from [163, 164, 165].
CLAS-data on $K\Sigma$ production by polarized beams initiated an updated large scale coupled-channels analysis of associated strange production on the nucleon. Based on the effective GiM Lagrangian, a combined CC analysis of $(\pi,\gamma)N\to K\Sigma$ hadro- and photo-production reactions was performed. The analysis covered a center of mass energy range up to $2~{}\mathrm{GeV}$. The central aim was to extract the resonance couplings to the $K\Sigma$ state. Both $s$-channel resonances and $t,~{}u$-channel background contributions are found to be important for an accurate description of angular distributions and polarization observables, assuring a high-quality description of the data. The extracted properties of isospin $I=3/2$ resonances were discussed in detail. In [160], it was found that the $I=1/2$ resonances are largely determined by the non-strangeness channels.
The calculations included 11 isospin $I=1/2$ resonances and 9 isospin $I=3/2$ resonances, respectively. The investigations were extended to the $I=1/2$ and $3/2$ sectors with the parameters fitted to newly published $K\Sigma$ photoproduction data together with the previous $\pi N\to K\Sigma$ measurements in the energy region $\sqrt{s}\leq 2.0~{}\mathrm{GeV}$. The included $K\Sigma$ photoproduction data are those of the $\gamma p\to K^{+}\Sigma^{0}$ published by the LEPS [166, 167, 168], CLAS [169, 170] and GRAAL [171] group, and those of $\gamma p\to K^{0}\Sigma^{+}$ released by the CLAS [172] and CBELSA [173] collaboration, respectively. The SAPHIR data have been left out here because of the known inconsistencies of the $K^{+}\Sigma^{0}$ data [161] with the corresponding CLAS and GRAAL data (for the details, see Ref. [170]). Also, the $K^{0}\Sigma^{+}$ SAPHIR data [161] have much bigger error bars than those of the CBELSA and CLAS groups. The data before 2002 are also no longer used. Results for total cross sections are shown in Fig. 26. Up to a total center-of-mass energy of about $\sqrt{s}=2~{}\mathrm{GeV}$, the data are well described.
Figure 26: Total cross sections for kaon production on the nucleon. Results of the Giessen model [174] are compared to CLAS, CBELSA, and SAPHIR data. Results of the full model calculation are shown in the left panel. Results using only the Born-amplitudes and t-channel meson exchange are displayed in the right panel.
The analysis included all charge channels, $K^{0}\Sigma^{\mp}$ and $K^{\pm}\Sigma^{0}$. A quite satisfactory description was achieved of the $\gamma p\to K^{+}\Sigma^{0}$ data ($\chi^{2}=1.8$) and the $\gamma p\to K^{0}\Sigma^{+}$ data ($\chi^{2}=2.0$). However, the pion-induced strangeness production reactions are described slightly less accurately as indicated by the corresponding $\chi^{2}$ values of $\chi^{2}=$ 4.1, 3.2, and 2.8 for the $\pi^{+}p\to K^{+}\Sigma^{+}$, $\pi^{-}p\to K^{0}\Sigma^{0}$, and $\pi^{-}p\to K^{+}\Sigma^{-}$ reactions, respectively. The parameters have been varied in our fit simultaneously to the $I=1/2$ and $3/2$ sectors. Although the new data are available with reduced total uncertainties, the refitted model parameters were changed only very little. A typical result is displayed in Fig. 27, illustrating the quality of the description on the example of $\pi^{-}p\to K^{0}\Sigma^{0}$ reaction. The complete set of results, including partial wave cross sections, angular distributions of cross sections, and polarization observables for the full set of $K\Sigma$ exit channels, is found in [174].
Figure 27: Differential cross section of $\pi^{-}p\to K^{0}\Sigma^{0}$ reaction. The solid (green), dashed (blue) and dotted (magenta) lines are the full model calculation, the model calculation with the $S_{11}(1650)$ and $F_{15}(1680)$ turned off, respectively. The numerical labels denote the center of mass energies in units of GeV. The figure is adapted from Ref. [174].
4 $\eta$-Meson Production
Understanding the dynamics of eta-meson production and, vice versa, the decay of nucleon resonances into the nucleon-eta exit channel is of ongoing interest in hadron spectroscopy. The $\eta$-meson photoproduction on the proton has been measured with high precision by the Crystal Ball Collaboration at MAMI [175]. These high-resolution data provide a new step forward in understanding the reaction dynamics and in the search for a signal from the “weak” resonance states. The main result reported in [175] is a very clean signal for a dip structure around $W=1.68~{}\mathrm{GeV}$, seemingly confirming older data [176, 177, 178, 179]. This raised the question of the origin of that structure, eventually indicating the appearance of a new narrow, possibly exotic, resonance state.
The aim of the study was to extend our previous coupled-channels analysis of the $\gamma p\to\eta p$ reaction by including the data from the new high-precision measurements [175]. The main question is whether the $\eta p$ reaction dynamics can be understood in terms of the established resonance states or whether a new state has to be introduced, thus confirming previous conjectures. A major issue for the analysis is unitarity and a consistent treatment of self-energy effects as visible in the total decay width of resonances. Since the latter are driven by hadronic interactions, the analysis of photo-production data requires the knowledge of the hadronic transition amplitudes as well. Hence, a coupled-channels description as in the Giessen model (GiM) is an indispensable necessity.
As discussed in every detail in Ref. [180] various relevant meson-baryon coupling constants were newly determined at the occasion of this work in large scale coupled-channels calculations. This gave rise to improved constraints on the interaction parameters and the derived resonance parameters, i.e., masses and widths. Representative examples are the mass and width of the $D_{13}(1520)$ resonance, $M=1516\pm 10~{}\mathrm{MeV}$ and $\Gamma=106\pm 4~{}\mathrm{MeV}$, agreeing with and confirming the values obtained earlier by Arndt et al. [181]. It is interesting to note that the mass of this resonance deduced from pion photoproduction tends to be $10~{}\mathrm{MeV}$ lower than the values derived from the pion-induced reactions [182]. The second $D_{13}(1900)$ state has a very large decay width. That state is likely to be related to the $D_{13}(2080)$ two-star resonance, proposed by PDG2024 [38].
The results of the calculation of the $\eta$-photo production channel are shown and compared with the experimental data in Fig. 28.
Figure 28: GiM results (full lines) of differential $\eta p$ cross section are compared to data from MAMI (symbols) data [175]. The figure is adapted from Ref. [183].
The calculations demonstrate a very satisfactory agreement with the experimental data in the whole kinematical region. The first peak is related to the $S_{11}(1535)$ resonance contribution. Similar to the $\pi^{-}p\to\eta n$ reaction, the $S_{11}(1650)$ and $S_{11}(1650)$ states interfere destructively, producing a dip around $W=1.68~{}\mathrm{GeV}$. The coherent sum of all partial waves leads to the more pronounced effect from the dip at forward angles.
Figure 29: GiM results for the photon-beam asymmetry $\Sigma$ in $\eta$ photoproduction on the proton compared to GRAAL data [179]. The figure is adapted from Ref. [180].
In Fig. 29, results for the photon-beam asymmetry $\Sigma$ are compared to the GRAAL data. One can see that even close to the $\eta N$ threshold where the calculations exhibit a dominant $S_{11}$ production mechanism, the beam asymmetry is non-vanishing for angles $\cos(\theta)\geq-0.2$. These results demonstrate that this observable is very sensitive to very small contributions from higher partial waves. At $W=1.68~{}\mathrm{GeV}$ and forward angles the GRAAL measurements show a rapid change of the asymmetry behavior. In [183] this effect was explained by a destructive interference between the $S_{11}(1535)$ and $S_{11}(1650)$ resonances which induces the dip at $W\simeq 1.68~{}\mathrm{GeV}$ in the $S_{11}$ partial wave. Note that the interference between $S_{11}(1535)$ and $S_{11}(1650)$ and the interference between different partial waves are of different nature. The overlapping of the $S_{11}(1535)$ and $S_{11}(1650)$ resonances does not simply mean a coherent sum of two independent contributions, but also includes rescattering (coupled-channel effects). Such interplay is hard to simulate by the simple sum of two Breit-Wigner distributions.
9 Double-Pion Production on the Nucleon
In certain energy regions, the $\pi N\to 2\pi N$ reaction accounts for up to 50% of the $\pi N$ inelasticity, as seen from Fig. 30. Therefore, this production channel must be included in any CC-PWA approach, as e.g., practiced in GiM. An improved and considerably extended description of double-pion production within our coupled channels scheme was started recently, and first results are found in Ref. [184]. The inclusion of multi-meson configurations into a coupled channels approach is a highly non-trivial exercise in three-body dynamics. In view of the complexities, physically meaningful approximations are necessary, retaining the essential dynamical aspects but making numerical calculations feasible. For that goal, the ansatz used in Ref. [184] relies on an isobar description of intermediate two-pion configurations and their decay into the final double-pion states on the mass shell. The derived processes contributing to the T-matrix of double-pion production on the nucleon in that energy region are depicted in Fig. 30.
This approach allows for the direct analysis of the $2\pi N$ experimental data. Since the corresponding Dalitz plots are found to be strongly non-uniform, it is natural to assume that the main effect to the reaction comes from the resonance decays into isobar sub-channels [185]. The most important contributions are expected to be from the intermediate $\sigma N$, $\pi\Delta$, and $\rho N$ states. The analysis of the $\pi N\to 2\pi N$ reaction would therefore provide very important information about the resonance decay modes into different isobar final states. The much richer baryon spectrum found in LQCD simulations [186, 187], the functional DSE approaches [188], and the CQM results [189, 190] that are observed in scattering experiments indicate the necessity for broader investigations, including a larger class of reaction channels. Experimentally, most of the non-strange baryonic states have been identified from the analysis of the elastic $\pi N$ data [181, 98]. As pointed out in [189] the signal of excited states with a small $\pi N$ coupling will be suppressed in the elastic $\pi N$ scattering. As a solution to this problem, a series of photoproduction experiments has been done to accumulate enough data for the study of the nucleon excitation spectra. However, the results from the photoproduction reactions are still controversial. While recent investigations of the photoproduction reactions presented by the BnGa group [191] reported indications for some new resonances, not all of these states are found in other calculations [182]. This raises a question about independent confirmation for the existence of such states from the investigations of other reactions.
Because of the smallness of the electromagnetic couplings, the largest contribution to the resonance self-energy comes from the hadronic decays. If the $N^{\ast}\to\pi N$ transition is small, one can expect sizable resonance contribution into remaining hadronic decay channels. As a result, the effect from the resonance with a small $\pi N$ coupling could still be significant in the inelastic pion-nucleon scattering: here the smallness of resonance coupling to the initial $\pi N$ states could be compensated by the potentially large decay branching ratio to other different inelastic final states. Such a scenario is realized, e.g., in the case of the well-known $N^{\ast}(1535)$ state. While the effect from this resonance to the elastic $\pi N$ scattering is only moderate at the level of the total cross section, its contribution to the $\pi N\to\eta N$ channel turns out to be dominant [180]. Since the $\pi N\to 2\pi N$ reaction could account for up to 50% of the total $\pi N$ inelasticity, this channel becomes very important not only for the investigation of the properties of already known resonances but also for the search for the signals of possibly unresolved states.
Figure 30: The processes contributing to double-pion production T-matrix are depicted diagrammatically: (a) and (b) production through the $\sigma$-isobar, (c) and (d) production through the $\Delta^{0}$-isobar. Symmetrization is indicated.
Another important issue in studies of the $2\pi N$ channel is related to the possibility to investigate cascade transitions like ${N^{\ast}}^{\prime}\to\pi N^{\ast}\to\pi\pi N$, where a massive state ${N^{\ast}}^{\prime}$ decays via intermediate excited $N^{\ast}$ or $\Delta^{\ast}$. It is interesting to check whether such decay modes are responsible for the large decay width of higher lying mass states. So far only the $\pi N^{\ast}(1440)$ isobar channel has been considered in a partial wave analysis (PWA) of the $\pi N\to 2\pi N$ experimental data [185].
There are several complications in the coupled-channel analysis of $2\to 3$ transitions. The first one is the difficulty to perform the partial-wave decomposition of the three-particle state. The second complication is related to the issue of three-body unitarity. For a full dynamical treatment of the $2\to 3$ reaction, the Faddeev equations (named after Ludwig Dmitrievich Faddeev (1934-2017)) have to be solved. Although appropriate theoretical and numerical methods are known, the effort inhibits practical implementations. Here, a coupled-channel approach for solving the $\pi N\to 2\pi N$ scattering problem in the isobar approximation is used, as widely practiced. In this formulation, the $(\pi/\pi\pi)N\to(\pi/\pi\pi)N$ coupled-channel equations are reduced to effective two-body scattering equations, taking advantage of intermediate two-body isobar production. Such a description accounts by construction for the full spectroscopic strength of intermediate channels and, in addition, provides a considerable numerical simplification.
Three-body unitarity leads to a relation between the imaginary part of the elastic scattering amplitude and the sum of the total elastic and inelastic cross sections by the well-known optical theorem. Since in the isobar approximation the pions in the $\pi\pi N$ channel are produced from the isobar sub-channels, all contributions to the total $\pi N\to\pi\pi N$ cross section are driven by the isobar production. The optical theorem can be fulfilled if all discontinuities in isobar sub-channels are taken into account. In the present work, the three-body unitarity is maintained up to interference terms between the isobar sub-channels.
The first resonance energy region is of particular interest because of the sizable effect from $N^{\ast}(1440)$. The dynamics of the Roper resonance turns out to be rich because of the two-pole structure reported in earlier studies [192, 193], (see [181, 194, 195] for the recent status of the problem). The origin of the Roper resonance is also controversial. For example, the calculations in the Jülich model explain this state as a dynamically generated pole due to the strong attraction in the $\sigma N$ sub-channel. At the same time, the Crystal Ball Collaboration finds no evidence of strong $t$-channel sigma-meson production in their $\pi^{0}\pi^{0}$ data [196]. From the further analysis of the $\pi^{0}\pi^{0}$ production, the effect of the sigma meson was found to be small [197]. On the other hand, the $pp\to pp\pi^{0}\pi^{0}$ scattering experiment by the CELSIUS-WASA collaboration [198] finds the $\sigma N$ decay mode of the Roper resonance to be dominant.
Figure 31: $\pi^{0}\pi^{0}$ differential cross sections for the reaction $\pi^{-}p\to\pi^{0}\pi^{0}n$ at fixed $t_{\pi^{0}\pi^{0}}=\cos{\theta_{\pi^{0}\pi^{0}}}$, shown in the upper left corner of the panels. Energy distributions for $-0.95\leq t_{\pi^{0}\pi^{0}}\leq+0.95$ are shown. The experimental data are from [197] The figure is adapted from Ref. [184].
In the region of the Roper resonance, our calculations are able to describe the mass distributions rather satisfactorily. Also, in this region, the production strength is shifted to higher invariant masses $m_{\pi^{0}\pi^{0}}^{2}$. At the same time, a peak at small $m_{\pi^{0}\pi^{0}}^{2}$ becomes also visible. In the present calculations, the fit tends to decrease the magnitude of the $\pi\Delta$ production and compensate it by enhancing the strength into $\sigma N$. The obtained decay branching ratio of $N^{\ast}(1440)$ for the $\sigma N$ channel is about twice as large as for the $\pi\Delta$.
Both the small peak at small and the broad structure at large invariant masses are well reproduced indicating an important interplay between the $\sigma N$ and $\pi\Delta$ production mechanisms. It is interesting that the isoscalar correlations in the $\pi\pi$ rescattering are also found to be necessary in order to reproduce the asymmetric shape of the mass distributions. Though the $\pi\Delta$ production gives rise to a two-peak structure only the first one at small $m_{\pi^{0}\pi^{0}}^{2}$ is visible at energies $1.4-1.468~{}\mathrm{GeV}$. Within the present calculation the second peak at high $m_{\pi^{0}\pi^{0}}^{2}$ is not seen because of the large $\sigma N$ contributions. In the present study $\pi^{0}\pi^{0}n$ production is calculated as a coherent sum of isobar contributions. Though the interference effects are important they are found to be very small at the level of the total cross sections.
To simplify the analysis, the $S_{11}$ and $P_{11}$ $\pi N$ partial waves are directly constrained by the single energy solutions (SES) derived by GWU(SAID) [181]. The experimental data on the $\pi^{-}p\to\pi^{0}\pi^{0}n$ reaction are taken from [197]. These measurements provide high-statistics data on the angular distributions ${\rm d}\sigma/{d\Omega_{\pi\pi}}$ where $\Omega_{\pi\pi}$ is the scattering angle of the $\pi\pi$ pair (or the final nucleon in c.m.).
The calculated $\pi^{0}\pi^{0}$ differential cross sections are shown in Fig. 31 and compared to the Crystal Ball data as a function of the c.m. energy. The measurements demonstrate a rapid rise of the cross sections at the energies $1.3-1.46~{}\mathrm{GeV}$, indicating the strong contribution coming from the Roper resonance, as also found in the GWU(SAID) [181] analysis.
The invariant $\pi^{0}\pi^{0}$ mass distributions play a crucial role in the separation of the isobar contributions. The $\pi^{-}p\to\pi^{0}\pi^{0}n$ reaction close to threshold is dominated by the $\sigma N$ production due to the $t-$ channel pion exchange. The nucleon Born term contribution to the $\pi\Delta$ channel is found to be less significant. The decay branching ratios of $N^{\ast}(1440)$ are obtained as $R_{\sigma N}^{N(1440)}=27^{+4}_{-9}\,\%$ and $R_{\pi\Delta}^{N(1440)}=12^{+5}_{-3}\,\%$.
The parameters extracted independently in the different approach of the BnGa group are $R_{\sigma N}^{N(1440)}=17^{+7}_{-7}\%$ and $R_{\pi\Delta}^{N(1440)}=21^{+8}_{-8}\%$, demonstrating that in spite of the visible difference in the central values, these quantities still coincide within error bars.
The double-pion channels are of high interest for $\chi$EFT. In [199], the Jülich group has been studying pion production off nucleons in the heavy baryon chiral perturbation theory to third order in the chiral expansion aiming at the determination of the low–energy constants. Most of the at that time available differential cross sections and angular correlation functions at low pion incident energies could be described together with total cross sections at higher energies. The contributions from the one-loop graphs were found to be essentially negligible once the dominant terms at second and third order are related to pion–nucleon scattering graphs with one pion were added. An interesting aspect is that the $\pi\pi N$ channels provide the possibility of extracting the pion–pion $S$–wave scattering lengths which otherwise are hard to access.
10 Double Pseudoscalar Meson Production Induced by Pions
1 Double-Pion Production by Pion Beams
Early studies of the dynamics of the $N\pi\pi$ final state were based on bubble chamber data on the reaction $\pi p\to\pi\pi N$ collected in non-polarized experiments by Berkeley, Saclay, and Rutherford laboratories (Fig. 32). 241,214 Bubble Chamber events have been analyzed in Isobar-model PWA at $W=1320-1930~{}\mathrm{MeV}$ by the Virginia Tech group led by Richard Arndt [185]. A summary of the number of Bubble Chamber events is given in Fig. 32.
Figure 32: Summary of the number of Bubble Chamber events for reactions $\pi p\to\pi\pi N$ analyzed at each energy.
Partial-wave inelasticities for elastic pion-nucleon scattering were determined with the aid of experimental data on $\pi N\to\pi\pi N$ processes in the beam-momentum range $300~{}\mathrm{MeV/c}<\mathrm{P_{beam}}<500~{}\mathrm{MeV/c}$ [200].
However, an obstacle for a much-desired joint PWA analysis is inconsistencies which at present inhibit merging $\pi N\to\pi N$ and $\pi N\to\pi\pi N$ databases. Problems exist specifically for small inelasticities (Fig. 33(left)), while the largest inelastic cross section related to the $P_{11}$ Roper resonance is in excellent agreement, being well described in SAID-SP06 for $\pi N\to\pi N$ [181] (Fig. 33(right)).
Figure 33: Partial-wave inelasticities $(1-\eta^{2})$ from the $\pi\to 2\pi$ analysis [200] (dotted curves) and $\pi N$ elastic scattering [181] (solid curves). Left: $S_{31}$, $P_{31}$, $P_{33}$, and $D_{33}$. Right: $P_{11}$.
A complete analysis of $\gamma N\to\pi\pi N$ ideally would require fitting all data obtained with both pion and photon beams. Already a long while ago, Pontecorvo considered a joint analysis of pion-nucleon scattering and single pion production in nucleon-nucleon and pion-nucleon interactions [201] but a full scale analysis of that kind is still pending.
2 Double-Pion Production and Polarization
An important set of observables on hadron structure and production dynamics is obtained in polarization measurements, see, e.g., the discussion around Fig. 29. For single pion production, such measurements are done at practically all hadron facilities and have become routine work in most of the CC approaches.
The situation is different, however, for double-pion production. The probably only double-pion production experiment including polarization measurements and subsequent theoretical analysis was done some time ago by Alekseev et al. [202] at ITEP (Moscow) . Data were taken for the reaction $\pi^{-}+\vec{p}\to\pi^{+}+\pi^{-}+p$ on polarized targets and at beam momentum of $1.78~{}\mathrm{Gev/c}$. By means of the SPIN spectrometer, designed especially for measurements of polarization observables in reactions leading to charged two- and three-particle final states, the full set of 14 spin observables could be measured. The primary aim of the experiment was to study specifically pion-pion interactions in the mass region of the $\rho(770)$ meson, but aiming at narrowing down the mass of the - until today - heavily disputed mass (and width) of the iso-scalar scalar $\sigma/f_{0}(500)$-meson (see Fig. 7 and the PDG2024 listings [38]).
The authors of Ref. [202] could indeed extract from the data highly valuable information on the $\pi^{+}\pi^{-}$ S-wave scattering phase shift in a PWA by inclusion of the polarization observables. However, the data obtained in that single measurement were not sufficient to determine unambiguously the desired mass parameter. Still, that - until today - single case experiment showed the important gain in information on the dynamics and spectroscopy of a two-body sub-system within the three-hadron final state, being populated in double-pion production on the nucleon.
3 Double Pseudoscalar Meson Production Beyond Pions
Double-kaon or eta meson production as well as vector meson (photo-)production are experimentally and theoretically highly demanding tasks. In a recent note, the ALICE collaboration announced first data on pion- and kaon-pair production [203]. From their mass spectra, they concluded that the states, which were observed in the mass region of $2~{}\mathrm{GeV}$, are well described by CQM calculations in the tradition of Karl and Isgur.
There is also ongoing research activity on the theoretical side. In [204] the authors presented a new approach describing double pion photoproduction off the nucleon in covariant chiral perturbation theory, thus connecting two important concepts of hadron physics.
The recent BnGa paper reported a combined analysis of photo and pion-induced double pion production [87]. Recent CBELSA/TAPS and CLAS (photo case) and Crystal Ball at BNL and HADES at CERN (pion case) collaboration data plus $\pi N$ PWA amplitudes from the SAID and Karlsruhe-Helsinki groups were involved in the analysis. The critical motivation of this study is the separation of $\pi\Delta$, $\rho N$, and $f^{0}(500)N$ final states because the experimental state for each case is the same, namely two pions and a nucleon. The CLAS experiment reported recently double-pion electroproduction on the proton and on deuterium [205], thus expanding the research into the domain of finite momentum transfer.
11 Hadron Production Induced by Nucleons, Nuclei, and Lepton Beams
Past measurements involving pion and kaon scattering measurements were made at a variety of laboratories, mainly in the 1970s and 1980s when experimental techniques were far inferior to the standards of today. In the US, pion beams in the momentum range $190~{}\mathrm{MeV/c}$ to $730~{}\mathrm{MeV/c}$ were available at the “Meson Factory” LAMPF in Los Alamos. This means that the maximum c.m. energy for baryon spectroscopy measurements at LAMPF was only $W\approx 1500~{}\mathrm{MeV}$. LAMPF was a linear accelerator for $1000~{}\mathrm{\mu A}$ of protons at $800~{}\mathrm{MeV}$. The meson factory PSI (former SIN) near Zurich was a sector-focused cyclotron capable of $100~{}\mathrm{\mu A}$ of protons at $600~{}\mathrm{MeV}$, and the Meson Factory TRIUMF in Vancouver was a sector-focused cyclotron for negative hydrogen ions up to $100~{}\mathrm{\mu A}$ at $500~{}\mathrm{MeV}$ [206, 207].
Important work is being done by experiments at other hadron facilities which were not mentioned in detail. Historically, SATURN at Saclay and the Synchro-Phasotron in Dubna, (JINR) for example, gave important insight into meson production processes on nuclei. The STAR@RHIC (BNL) and the GSI at Darmstadt experiments produced - and are producing - a wealth of data on meson and kaon, and partially also vector meson production in heavy ion collisions at relativistic energies. The meson yields of such processes, occurring in compressed and/or heated nuclear matter, were essential for understanding the dynamics of stressed baryon matter. Practically all currently operating LHC experiments are searching for exotic hadrons and nuclei, exploring, e.g., flavored nuclei and antimatter-nuclei, the latter used for searching signals of CP and CPT violation. COMPASS was and AMBER is devoted to meson physics, both located at CERN. J-PARC at Tokai has a strong program on hadron spectroscopy. An own class of facilities are the electron-positron colliders BESIII and BELLE. The experiments are mainly focused on flavor physics.
1 Resonance Spectrum of the Nucleon
Coupled channels approaches like the ones discussed in the previous sections are a valuable tool for understanding the spectrum of the excited nucleon. A natural question is how well the spectra of CC approaches agree with the results of other independent spectroscopic investigations?
Here, we address that issue for the case of the GiM coupled channels approach which may serve as a representative example. As mentioned before, the GiM is based on a phenomenological covariant field theory of baryons, mesons, and their interactions. By construction, the fundamental symmetries of QCD, including also chiral symmetry, are conserved. Resonance and background contributions are generated consistently out of the tree-level interactions as defined by the underlying Lagrangian.
The scattering amplitudes are determined by a linear system of coupled equations which is solved numerically in K-matrix approximation and partial wave representation. The gauge-invariant description of high-spin resonances was discussed in detail. Applications to selected reaction channels have been presented, ranging from single pion, eta, and kaon production to double-pion production to investigations of omega and $K^{*}$ vector meson channels.
Hence, CC approaches like the GiM and the other aforementioned projects incorporate the defining elements of hadron physics on the theoretical level. The approaches use comparable numerical methods for solving the CC problem where attention is paid to the fact that the numerical procedures conserve the basic symmetries and conservation laws.
The GiM baryon level scheme resulting from the investigations of meson production channels is summarized in Fig. 34 and compared to the observed spectrum as found in the PDG resonance compilations [182, 38]. It is noteworthy that the GiM spectrum is the result of a network of calculations in which resonances of one kind serve as intermediate states in analyses of resonances in reactions of another kind, e.g., the $\pi\Delta$ channels in double-pion production. The rather satisfactory agreement is a strong and encouraging confirmation of CC approaches as a tool for spectroscopic research in hadron production reactions.
Figure 34: Comparison of the GiM $N^{\ast}$ resonance level scheme (red bars) in various partial waves to the PDG compilation [182] (blue bars).
12 Model-Independent Analysis and Optimized Storage of Hadronic Reaction Data
The previous section was devoted to an involved multi-channel description of experimental data. For that purpose, models about the reaction mechanism and the interactions had to be invented, mostly on phenomenological grounds. Hence, model parameters were adjusted by fits to data. As seen, those approaches are successful in the sense of reproducing data in a consistent manner which, in view of the complexity of the task, is a remarkable achievement.
However, modern accelerator facilities in combination with detectors are capable of providing a tremendous amount of experimental data. Here the problem arises, how to present those numerous detailed data in a model-independent manner and make them available - and manageable - for further investigations. The standard approach is graphical representation, displaying various cuts through the multi-dimensional volume of measured data resulting in a multitude of graphs. That approach is necessarily selective and lacking complete coverage.
As an alternative, Azimov et al. [208] has suggested expanding experimental data into a series of a complete set of orthogonal functions, e.g., Legendre polynomials $P_{J}(z)$. For unpolarized differential cross sections, measured angular distributions, for example, may be decomposed as
$\frac{d\sigma}{dz}(W,z)=\sum_{J=0}^{\infty}A^{(\sigma)}_{J}(W)P_{J}(z)\>,$ (24)
where where $W$ is the c.m. energy, $z=\cos\theta$, and $\theta$ is the polar c.m. angle. Formally, this series is infinite. However, in real situations only a finite number of Legendre coefficients $A^{(\sigma)}_{J}(W)$ are effectively necessary. As a cut-off criterion one may impose that only those are considered with coefficients whose central values are larger by a given limit than the estimated error. The method was illustrated, for instance, in Ref. [209] using photoproduction data of the reaction $\gamma p\to\pi^{0}p$. As a result, the whole set of data appears to be represented in a compact form of energy dependent Legendre coefficients (Fig. 35), easy to use in later theoretical re-analyses or any other application of the data. The approach may be optimized to other kinds of data by choosing function systems well adapted to the case. Such compression strategies are used widely in modern data management and data mining.
Figure 35: Top panel: Samples of the $\gamma p\to\pi^{0}p$ differential cross sections, $d\sigma/d\Omega$, from A2 Collaboration at MAMI measurements (blue filled circles) [209] with the best fit results using Legendre polynomials (red dashed lines). The error bars on all data points represent statistical uncertainties only. Values of $E$ in each plot indicate the laboratory photon energies. Bottom panel: Coefficients of Legendre polynomials (blue filled circles). The error bars of all values represent $A^{(\sigma)}_{J}(W)$ uncertainties from the fits in which only the statistical uncertainties were used. Solid lines are plotted to help guide the eye. Red vertical arrows indicate masses of the four-star resonances (BW masses) known in this energy range [38]. The upper row of arrows corresponds to $N^{\ast}$ states with isospin $I=1/2$ and the lower row corresponds to $\Delta^{\ast}$ with $I=3/2$. Adopted from Ref. [208].
The Legendre coefficients have also other, less evident properties. For parity conserving reactions, some of those properties are related to partial-wave amplitudes for states of definite parity. In the unpolarized cross section the positive- and negative-parity amplitudes always appear symmetrically (that is why the unpolarized cross section by itself does not allow to determine the parity of a particular state). It is not quite so for $A^{(\sigma)}_{J}$. One can show that the Legendre coefficients provide specific discrimination of parities: $A^{(\sigma)}_{J}$ with odd $J$ contain only interferences of states with opposite parities, while $A^{(\sigma)}_{J}$ with even $J$ contain only interferences of states with the same parities, positive or negative. And, of course, only the even $J$ coefficients may contain squares of absolute values of various partial-wave amplitudes.
One can summarize the above statements in the following way:
  • Legendre expansions provide a model-independent approach suitable for presentation of modern detailed (high-precision and high-statistics) data for two-hadron reactions.
  • This approach is applicable both to cross sections and to polarization variables; it is much more compact than traditional methods, at least, at not very high energies.
  • The Legendre coefficients reveal specific correlations and interferences between states of definite parities.
  • Due to interference with resonances, high momentum Legendre coefficients open a window to study higher partial-wave amplitudes, which are out-of-reach in any other ways.
Concluding this, one should emphasize that direct interference has become a useful instrument to search for and study rare decays of well-established resonances. However, its possibilities are limited by restrictions for the resonance quantum numbers. Rescattering interference is not limited by such requirements and, therefore, may provide effective methods to search for and study new resonances with arbitrary quantum numbers. Data on multi-hadron decays of heavy particles also present a new rapidly expanding area for applications of different kinds of interference both to study the spectroscopy of resonances and to establish their characteristics.
13 Summary and Outlook
The long journey from first, at that time not understood signals observed by Victor in high-altitude balloon campaigns, and in the following years increasingly being confirmed in different independent experiments to modern hadron physics was reviewed. The improving experimental techniques were accompanied by epochal developments in theory, changing forever our understanding of the material world and the origin of matter. That exciting journey is by far not coming to an end - it is ongoing and many surprises can be expected for the future.
The final goal - yet waiting to be achieved - is to connect the results derived by CC methods from data to QCD predictions as obtained by LQCD and the respective functional approaches discussed in this Encyclopedia. The iterative approach, eventually converging to a (self-) consistent picture, is indicated in Fig. 36.
Figure 36: Illustration of the long-term goal of hadron physics, aiming at connecting hadron phenomenology to QCD.
Acknowledgments
This work was supported in part by the U. S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award No. DE–SC0016583 (I.S.) and Deutsche Forschungsgemeinschaft (DFG), grants Le439/16 and Le439/17 (H.L.).